### v1

# Electroweak Monopole-Antimonopole Pair Production at LHC

Abstract
One of the urgent issues in high energy physics is the experimental confirmation of the electroweak monopole predicted by the standard model, and currently MoEDAL at LHC is actively searching for the monopole. However, the present LHC cannot produce the monopole if the mass is bigger than 7 TeV, while the monopole mass is expected to be around $M_{W}/\alpha\simeq 11~{}\text{TeV}$. In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass is bigger than 7 TeV, based on the following ideas. First, in the topological production of the monopole the baby monopole mass at creation could be considerably smaller than the adolescent mass. Second, the binding energy of the monopole-antimonopole pair could effectively reduce the mass of the bound state. We discuss how these ideas can actually be realized at LHC to produce the monopole pairs. In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.2 TeV, smaller than the adolescent monopole mass around 11.0 TeV. Moreover, we show that LHC could produce the monopolium bound state with mass around 5.3 TeV, even when the total mass of the monopole-antimonopole pair is around 22.0 TeV. Our analysis could play an important role for MoEDAL experiment.
electroweak monopole, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
I Introduction
With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2, 3]. After the Dirac monopole we have had the Wu-Yang monopole, the ’t Hooft-Polyakov monopole, the grand unification monopole [4, 5, 6, 7, 8, 9, 10, 11]. But the electroweak monopole (also known as the Cho-Maison monopole), despite the initial skepticism on the existence of such monopole in the standard model [12, 13], stands out as the most realistic monopole that could exist in nature and could actually be detected [14, 15, 16].
Indeed the Dirac monopole in electrodynamics should transform to the electroweak monopole after the unification of the electromagnetic and weak interactions, and the Wu-Yang monopole in QCD is supposed to make the monopole condensation to confine the color. Moreover, the ’tHooft-Polyakov monopole exists only in an hypothetical theory, and the grand unification monopole which could have been amply produced at the grand unification scale in the early universe probably has become completely irrelevant at present universe after the inflation [17, 18].
This makes the experimental confirmation of the electroweak monopole one of the most urgent issues in the standard model after the discovery of the Higgs particle. In fact the detection of this monopole, not the Higgs particle, should be regarded as the final (and topological) test of the standard model. For this reason the newest MoEDAL detector at LHC is actively searching for the monopole [19, 20, 21, 22].
To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [14, 15, 23, 24, 25, 26, 27, 28]. First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’t Hooft-Polyakov monopoles. Second, the magnetic charge of the monopole is not $2\pi/e$ but $4\pi/e$, twice that of the Dirac monopole. This is because the period of the electromagnetic U(1) subgroup of the standard model is $4\pi$, since the electromagnetic U(1) (partly) comes from the U(1) subgroup of SU(2). Third, the mass of the monopole $M$ is of the order of several TeV, probably between 4 to 10 TeV. This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the coupling is magnetic (i.e., $4\pi/e$). This makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV. Despite this, the size of the monopole is set by the weak boson masses. This is because the monopole solution has the Higgs and W boson dressing which has the exponential damping fixed by the weak boson masses.
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass coud turn out to be bigger than this, around 11.0 TeV. If this is the case, we may have little chance to detect the monopole at LHC. In this case we may have to try to detect the remnant monopoles at present universe produced in the early universe [27], or wait for the next upgrade of LHC.
However, two observations could make the monopole signal at LHC non-zero even when the mass becomes over 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole, or the quantized magnetic flux of the monopole-antimonopole pair. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass is bigger than 7 TeV.
The second point is related to the electroweak monopole production mechanism at LHC. At present there are three contending monopole production mechanisms. Two popular ones in the high energy physics community are the Drell-Yan (and two photon fusion) process [19, 20] and the Schwinger mechanism [22, 29, 30]. The third one is the topological monopole production mechanism, in which the thermal fluctuation of the Higgs vacuum which takes place during the electroweak phase transition produces the monopole [27]. LHC could produce the monopole with this mechanism because the fireball of the p-p collision could re-create the electroweak phase transition in the early universe, and thus produces the monopoles topologically by the thermal fluctuation of the Higgs vacuum.
A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass is heavier than 7 TeV. To understand this point we first have to keep in mind that the monopole mass we have been talking about is the mass at the zero temperature. However, in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [27]. This tells that we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV. We show that in the topological thermal production the monopole production at LHC starts from $T_{c}\simeq 147~{}\text{TeV}$ and stops at $T_{min}^{LHC}\simeq 119~{}\text{TeV}$ before the Ginzburg temperature $T_{G}\simeq 57.5~{}\text{GeV}$, with the monopole mass around 5.3 TeV. Moreover, we show that LHC could produce the monopolium bound states of mass around 5.3 TeV, even though the mass of the monopole-antimonopole pair becomes 22.0 TeV.
As we will see, just after the phase transition the baby monopole mass could be as low as 1.0 TeV. Notice that this would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist. On the other hand, we emphasize that the monopole production in the early universe and the monopole production at LHC have important differences. First, the monopole production in the early universe stops around the Ginzburg temperature, while the monopole production at LHC stops at higher temperature. This is because at the present LHC the monopole mass can not exceed 7 TeV. So LHC has less monopole production time. Second, the size of the thermal radiation bath in the early universe is limited only by the horizon distance, but the size of the fireball at LHC in p-p collision is fixed by the proton size which is very small. Because of this, the monopole in the early universe has unlimited energy nearby but the monopole at LHC has very limited energy coming from the p-p fireball. These differences play important roles in the monopole production at LHC.
The existence of a monopole-antimonopole bound state, the metastable quantized magnetic flux made of Cho-Maison monopole-antimonopole pair, has recently been confirmed [31]. It has the expected pole separation length around $8.2/M_{W}$ and has the magnetic dipole moment roughly $8.6/M_{W}$. This extended size (length) of the magnetic flux suggests that the energy of the pair might be bigger than the sum of the monopole and antimonopole masses. This implies that LHC might need more energy to produce the magnetic flux made of the monopole-antimonopole pair than to produce the monopole and antimonopole separately.
This makes the monopolium bound state another probable signal for the monopole production at LHC. The monopolium itself has been studied before for various reasons [32, 33, 34]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV, the mass scale of the low lying hadrons [34]. But the electroweak monopolium made of the Cho-Maison monopole pair has not been studied before.
Our study on the electroweak monopolium tells the followings. First, the naive atomic model of monopolium does not describe a realistic monopolium, because it has unrealistically big binding energy and too small size. This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is huge.
But we could avoid these difficulties and have a realistic electroweak monopolium modifying the magnetic potential at short distance. AS we will see, in this modification we can show that the binding energy reduces the monopolium mass from 22.0 TeV to 5.3 TeV. This is precisely what we hoped for. With this we could conclude that LHC could produce the electroweak monopolium of the mass considerably lighter than 14 TeV without much difficulty. In this case we could try to detect the monopolium bound state experimentally detecting the decay modes.
The paper is organized as follows. In Section II we discuss the electroweak monopole production mechanism at LHC. In Section III we review the monopole production in the electroweak phase transition in the early universe. In Section IV we discuss how LHC could produce the monopole in a similar manner, and argue that LHC could produce the baby electroweak monopole whose mass is smaller than the adolescent mass. In Section V we discuss the naïve Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium. In Section VI we discuss a more realistic electroweak monopolium that could be produced at LHC. In Section VII we discuss the physical implications of our results.
II Electroweak Monopole Production Mechanisms at LHC
Before we discuss the electroweak monopole production at LHC, we review the standard model. Let us start from the (bosonic sector of the) Weinberg-Salam Lagrangian,
${\cal L}_{WS}=-|{\cal D}_{\mu}\phi|^{2}-\frac{\lambda}{2}\big{(}\phi^{\dagger}\phi-\frac{\mu^{2}}{\lambda}\big{)}^{2}-\frac{1}{4}\vec{F}_{\mu\nu}^{2}-\frac{1}{4}G_{\mu\nu}^{2},$
${\cal D}_{\mu}\phi=\big{(}\partial_{\mu}-i\frac{g}{2}\vec{\tau}\cdot{\vec{A}}_{\mu}-i\frac{g^{\prime}}{2}B_{\mu}\big{)}\phi,$ (1)
where $\phi$ is the Higgs doublet, $\vec{F}_{\mu\nu}$ and $G_{\mu\nu}$ are the gauge field strengths of $SU(2)$ and $U(1)_{Y}$ with the potentials ${\vec{A}}_{\mu}$ and $B_{\mu}$, and $g$ and $g^{\prime}$ are the corresponding coupling constants, and $\cal D_{\mu}$ is the covariant derivative of the $SU(2)$ and $U(1)_{Y}$ gauge groups. Expressing $\phi$ with the Higgs field $\rho$ and unit doublet $\xi$ by
$\phi=\dfrac{1}{\sqrt{2}}\rho~{}\xi,~{}~{}~{}(\xi^{\dagger}\xi=1),$ (2)
we have
${\cal L}_{WS}=-\frac{1}{2}(\partial_{\mu}\rho)^{2}-\frac{\rho^{2}}{2}|{\cal D}_{\mu}\xi|^{2}-\frac{\lambda}{8}\big{(}\rho^{2}-\rho_{0}^{2}\big{)}^{2}$
$-\frac{1}{4}\vec{F}_{\mu\nu}^{2}-\frac{1}{4}G_{\mu\nu}^{2},$ (3)
where $\rho_{0}=\sqrt{2\mu^{2}/\lambda}$ is the vacuum expectation value of the Higgs field.
We can express (1) in terms of the physical fields gauge independently [27]. With the Abelian decomposition of ${\vec{A}}_{\mu}$
${\vec{A}}_{\mu}={\hat{A}}_{\mu}+{\vec{W}}_{\mu},$
${\hat{A}}_{\mu}=A_{\mu}{\hat{n}}-\frac{1}{g}{\hat{n}}\times\partial_{\mu}{\hat{n}},~{}~{}~{}{\hat{n}}=-\xi^{\dagger}\vec{\tau}\xi,$
${\vec{F}}_{\mu\nu}={\hat{F}}_{\mu\nu}+{\hat{D}}_{\mu}{\vec{W}}_{\nu}-{\hat{D}}_{\nu}{\vec{W}}_{\mu}+g{\vec{W}}_{\mu}\times{\vec{W}}_{\nu},$
${\hat{D}}_{\mu}=\partial_{\mu}+g{\hat{A}}_{\mu}\times,$ (4)
we have
${\vec{F}}_{\mu\nu}^{2}={F^{\prime}}_{\mu\nu}^{2}+2|D^{\prime}_{\mu}W_{\nu}-D^{\prime}_{\nu}W_{\mu}|^{2}-4igF^{\prime}_{\mu\nu}W_{\mu}^{*}W_{\nu}$
$-g^{2}(W_{\mu}^{*}W_{\nu}-W_{\nu}^{*}W_{\mu})^{2},$
$F^{\prime}_{\mu\nu}=\partial_{\mu}A^{\prime}_{\nu}-\partial_{\nu}A^{\prime}_{\mu},$
$A^{\prime}_{\mu}=A_{\mu}+C_{\mu},~{}~{}~{}C_{\mu}=-\frac{2i}{g}\xi^{\dagger}\partial_{\mu}\xi$
$D^{\prime}_{\mu}=\partial_{\mu}+igA^{\prime}_{\mu},~{}~{}~{}W_{\mu}=\frac{W^{1}_{\mu}+iW^{2}_{\mu}}{\sqrt{2}}.$ (5)
With this we can define $A^{\rm(em)}$ and $Z_{\mu}$ by
$\left(\begin{array}[]{cc}A_{\mu}^{\rm(em)}\\ Z_{\mu}\end{array}\right)=\frac{1}{\sqrt{g^{2}+g^{\prime 2}}}\left(\begin{array}[]{cc}g&g^{\prime}\\ -g^{\prime}&g\end{array}\right)\left(\begin{array}[]{cc}B_{\mu}\\ A^{\prime}_{\mu}\end{array}\right)$ (12)
$=\left(\begin{array}[]{cc}\cos\theta_{\rm w}&\sin\theta_{\rm w}\\ -\sin\theta_{\rm w}&\cos\theta_{\rm w}\end{array}\right)\left(\begin{array}[]{cc}B_{\mu}\\ A_{\mu}^{\prime}\end{array}\right),$ (17)
and have the the identity
$|{\cal D}_{\mu}\xi|^{2}=\frac{1}{4}Z_{\mu}^{2}+\frac{g^{2}}{2}W_{\mu}^{*}W_{\mu}.$ (18)
From this we can express the Weinberg-Salam Lagrangian in terms of the physical fields,
${\cal L}_{WS}=-\frac{1}{2}(\partial_{\mu}\rho)^{2}-\frac{\lambda}{8}\big{(}\rho^{2}-\rho_{0}^{2}\big{)}^{2}$
$-\frac{1}{4}{F_{\mu\nu}^{\rm(em)}}^{2}-\frac{1}{4}Z_{\mu\nu}^{2}-\frac{g^{2}}{4}\rho^{2}W_{\mu}^{*}W_{\mu}-\frac{g^{2}+g^{\prime 2}}{8}\rho^{2}Z_{\mu}^{2}$
$-\frac{1}{2}|(D_{\mu}^{\rm(em)}W_{\nu}-D_{\nu}^{\rm(em)}W_{\mu})+ie\frac{g}{g^{\prime}}(Z_{\mu}W_{\nu}-Z_{\nu}W_{\mu})|^{2}$
$+ieF_{\mu\nu}^{\rm(em)}W_{\mu}^{*}W_{\nu}+ie\frac{g}{g^{\prime}}Z_{\mu\nu}W_{\mu}^{*}W_{\nu}$
$+\frac{g^{2}}{4}(W_{\mu}^{*}W_{\nu}-W_{\nu}^{*}W_{\mu})^{2},$ (19)
where $D_{\mu}^{\rm(em)}=\partial_{\mu}+ieA_{\mu}^{\rm(em)}$ and $e$ is the electric charge
$e=\frac{gg^{\prime}}{\sqrt{g^{2}+g^{\prime 2}}}=g\sin\theta_{\rm w}=g^{\prime}\cos\theta_{\rm w}.$ (20)
We emphasize that this is not the Weinberg-Salam Lagrangian in the unitary gauge. This is the gauge independent abelianization of the Weinberg-Salam Lagrangian. Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson masses are given by $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [14]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times heavier than the W boson mass, around 11 TeV. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$). We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [23, 24, 25, 27, 28]. In the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
Figure 1: The Feynman diagrams of the popular monopole production mechanism at LHC given by Drell-Yan and two-photon fusion process. Here the proton pair could also be interpreted as the quark pair.
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism. There are two contending monopole production mechanisms in particle physics community at present, the Drell-Yan (and two photon fusion) process and the Schwinger mechanism. In the Drell-Yan process the monopole-antimonopole pairs are produced by the electromagnetic interaction through the photon [19, 20]. But in the Schwinger mechanism a strong magnetic field creates the monopole-antimonopole pairs, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [22, 29, 30]. The justification for this comes from the fundamental assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ with $(e,g)$ to $(g,-e)$. The Drell-Yan process is shown graphically in Fig. 1.
Notice, however, that something is strange with these pictures. For instance, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process. But the Weinberg-Salam Lagrangian has no interaction which can be described by the above Feynman diagram. Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background. Most importantly, the monopole is not an ordinary point particle but a topological particle, so that we need a change of topology to produce it. But it is not clear how the Drell-Yan process or the Schwinger mechanism could accommodate this topological nature of the monopole.
Figure 2: The topological monopole production mechanism in the early universe or at LHC induced by the thermal fluctuation of the Higgs vacuum.
The standard model provides another monopole production mechanism, the topological monopole production generated by the thermal fluctuation of the Higgs vacuum which induces the change of topology, which is the electroweak monopole production mechanism in the early universe [27]. In this picture the monopoles are produced in the course of the electroweak phase transition by thermal fluctuations of the Higgs vacuum, which takes place after the phase transition and stops at the Ginzburg temperature [27]. This is shown in Fig. 2, which should be contrasted with Fig. 1.
A unique feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass at creation (the baby monopole mass) could be considerably smaller than the adolescent monopole mass. This is because in this thermal production, the monopole mass depends on the temperature at creation. As we will see in the following this could have a deep impact on the monopole production at LHC.
If so, what is the the monopole production mechanism at LHC? We do not know yet. On the other hand, it has often been claimed that LHC should reproduce the early universe. If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production the early universe. In this paper we will adopt this view and discuss how LHC could produce the electroweak monopole. For this we have to understand the electroweak monopole production in the early universe.
III Monopole Production in Early Universe: A Review
It has generally been believed that the monopole production in the early universe critically depends on the type of the phase transition. In the first order phase transition the vacuum bubble collisions in the unstable vacuum are supposed to create the monopoles through the quantum tunneling to the stable vacuum during the phase transition, so that the monopole production is supposed to be suppressed exponentially by the vacuum tunneling [17]. On the other hand the monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism with no exponential suppression [35, 36].
We emphasize, however, that this popular view may have a critical defect [27]. This is because in the topological monopole production mechanism the thermal fluctuation of the Higgs vacuum which provides the seed of the monopoles continues untill the temperature drops to the Ginzburg temperature [37]. This means that, even in the first order phase transition we can have the monopole production after the vacuum bubble tunneling without the exponential suppression, if the Ginzburg temperature becomes less than the critical temperature. In this case the monopole production in the first order phase transition becomes qualitatively the same as in the second order phase transition. This tells that the popular exponential suppression of the monopole production in the first order phase transition is only half of the full story which could be totally misleading.
In the second order phase transition the thermal fluctuation of the Higgs vacuum could also modify the Kibble-Zurek mechanism considerably, because it provides more time for the monopole production as the thermal fluctuation could continue long after the phase transition. This tells that what is important in the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
To amplify this point we start from the temperature dependent effective action of the standard model expressed by [27, 38, 39, 40, 41, 42, 43]
$V_{eff}(\rho)=V_{0}(\rho)-\frac{C_{1}}{12\pi}\rho^{3}~{}T+\frac{C_{2}}{2}\rho^{2}~{}T^{2}-\frac{\pi^{2}}{90}N_{*}T^{4}$
$+\delta V_{T},$
$V_{0}(\rho)=\frac{\lambda}{8}(\rho^{2}-\rho_{0}^{2})^{2},$
$C_{1}=\frac{6M_{W}^{3}+3M_{Z}^{3}}{\rho_{0}^{3}}\simeq 0.36,$
$C_{2}=\frac{4M_{W}^{2}+2M_{Z}^{2}+M_{H}^{2}+4m_{t}^{2}}{8\rho_{0}^{2}}\simeq 0.37,$ (21)
where $V_{0}$ (with $\lambda\simeq 0.26$ and $\rho_{0}\simeq 246~{}{\rm GeV}$) is the zero-temperature potential, $N_{*}$ is the total number of distinct helicity states of the particles with mass smaller than $T$ (counting fermions with the factor 7/8), $C_{1}$ and $C_{2}$ are the contributions from the $\rho$ dependent part of the masses of gauge bosons, Higgs field, and fermions, $m_{t}$ is the top quark mass, and $\delta V_{T}$ is the slow-varying logarithmic corrections and the lighter quark contributions which we will neglect from now on.
The potential has three local extrema at
$\rho_{s}=0,$
$\rho_{\pm}(T)=\Big{\{}\frac{C_{1}}{4\pi\lambda}\pm\sqrt{\Big{(}\frac{C_{1}}{4\pi\lambda}\Big{)}^{2}+\frac{\rho_{0}^{2}}{T^{2}}-\frac{2C_{2}}{\lambda}}\Big{\}}~{}T^{.}$ (22)
The first extremum $\rho_{s}=0$ represents the Higgs vacuum of the symmetric (unbroken) phase, the second one $\rho_{-}(T)$ represents the local maximum, and the third one $\rho_{+}(T)$ represent the local minimum Higgs vacuum of the broken phase. But the two extrema $\rho_{\pm}$ appear only when $T$ becomes smaller than $T_{1}$
$T_{1}=\frac{4\pi\lambda}{\sqrt{32\pi^{2}\lambda C_{2}-C_{1}^{2}}}~{}\rho_{0}\simeq 146.74~{}\text{GeV}.$ (23)
So above this temperature only $\rho_{s}=0$ becomes the true vacuum of the effective potential, and the electroweak symmetry remains unbroken.
At $T=T_{1}$ we have
$\rho_{-}=\rho_{+}=(C_{1}/4\pi\lambda)~{}T_{1}\simeq 16.3~{}{\rm GeV},$ (24)
but as temperature cools down below $T_{1}$ we have two local minima at $\rho_{s}$ and $\rho_{+}$ with $V_{eff}(0)<V_{eff}(\rho_{+})$, until $T$ reaches the critical temperature $T_{c}$ where $V_{eff}(0)$ becomes equal to $V_{eff}(\rho_{+})$,
$T_{c}=\sqrt{\frac{18}{36\pi^{2}\lambda C_{2}-C_{1}^{2}}}~{}\pi\lambda\rho_{0}\simeq 146.70~{}{\rm GeV},$
$\rho_{+}(T_{c})=\frac{C_{1}}{3\pi\lambda}~{}T_{c}\simeq 21.7~{}{\rm GeV}.$ (25)
So $\rho_{s}=0$ remains the minimum of the effective potential for $T>T_{c}$. Notice that $\rho_{+}(T_{c})/\rho_{0}\simeq 0.09$.
At $T_{c}$ the new vacuum bubbles start to nucleate at $\rho_{s}=0$, which tunnels to the stable vacuum $\rho_{+}$ after $T_{c}$. Below this critical temperature $\rho_{+}$ becomes the true minimum of the effective potential, but $\rho_{s}=0$ remains a local (unstable) minimum till the temperature reaches $T_{2}$,
$T_{2}=\sqrt{\frac{\lambda}{2C_{2}}}~{}\rho_{0}\simeq 146.42~{}\text{GeV},$
$\rho_{+}(T_{2})=\frac{C_{1}}{2\pi\lambda}~{}T_{2}\simeq 32.5~{}\text{GeV}.$ (26)
From this point $\rho_{+}$ becomes the only (true) minimum, which approaches to the well-known Higgs vacuum $\rho_{0}$ at zero temperature. The effective potential (21) is shown in Fig. 3.
This tells that the electroweak phase transition is of the first order. However, notice that the energy barrier is extremely small,
$\frac{V_{eff}(\rho_{-})-V_{eff}(\rho_{+})}{V_{eff}(\rho_{+})}\Big{|}_{T_{c}}\simeq 3.8\times 10^{-6}.$ (27)
Moreover, the barrier lasts only for short period since the temperature difference from $T_{1}$ to $T_{c}$ is very small, $\delta=(T_{1}-T_{c})/T_{c}\simeq 0.0002$. So for all practical purposes we could treat the electroweak phase transition as a second order phase transition.
Figure 3: The effective potential (21) at different temperatures. Notice that the potentials at $T_{1},~{}T_{c},~{}T_{2}$ are almost indistinguishable. Here the unit of $V_{eff}$ is chosen to be $V_{0}=(\lambda/8)\rho_{0}^{4}=1$.
The monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism, so that the monopole production start from $T_{c}$. However, we have to keep in mind that the thermal fluctuations of the Higgs vacuum which create the seed of the monopoles continue as long as [27, 37]
$\xi^{3}\Delta F\leq T,~{}~{}~{}~{}\Delta F(T)=V(\rho_{s})-V(\rho_{+}),$ (28)
where $\xi(T)$ is the correlation length of the Higgs field and $\Delta F(T)$ is the difference in free energy density between two phases. This large fluctuation disappears when the equality holds, at the Ginzburg temperature $T_{G}$.
We can find it graphically from (21) and (28). This is shown in Fig. 4. From this we have [27]
$T_{G}\simeq 57.49~{}\text{GeV},~{}~{}~{}~{}\rho_{+}(T_{G})\simeq 232.9~{}\text{GeV}.$ (29)
The effective potential at the Ginzburg temperature is shown in Fig. 3. Notice that at $\rho_{+}(T_{G})$ is quite close to the Higgs vacuum at zero temperature, which confirms that the electroweak monopole production lasts long time after the phase transition.
Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the black line represents $T_{G}$.
With this observation we can say that the monopole formation takes place between $T_{c}$ and $T_{G}$, or roughly around $T_{i}$,
$T_{i}=\frac{T_{c}+T_{G}}{2}\simeq 102.1~{}\text{GeV},$
$\rho_{+}(T_{i})\simeq 188.4~{}\text{GeV}.$ (30)
To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [44]
$t=\Big{(}\frac{90}{32\pi GN_{*}(T)}\Big{)}^{1/2}~{}\frac{1}{T^{2}}.$ (31)
So, with $N_{*}\simeq 385$ (including $\gamma,\nu,g,e,\mu,\pi,u,d,s,c,b,\tau,\\ W,Z,H$) we have
$t\simeq 0.048\times\frac{M_{P}}{T^{2}}\simeq 3.9\times 10^{-7}\big{(}\frac{\rm GeV}{T}\big{)}^{2}sec.$ (32)
From this we can say that the electroweak monopole production start from $1.8\times 10^{-11}sec$ to $1.2\times 10^{-10}sec$ after the big bang for the period of $10.3\times 10^{-11}~{}sec$, or around $3.5\times 10^{-11}sec$ after the big bang in average.
The effective potential (21) gives us two important parameters of the electroweak phase transition, the temperature dependent Higgs mass $\bar{M}_{H}$ which determine the correlation length $\xi=1/\bar{M}_{H}$,
$\bar{M}_{H}^{2}=\frac{d^{2}V_{eff}}{d\rho^{2}}\Big{|}_{\rho_{min}}$
=$=\left\{\begin{array}[]{ll}\dfrac{1}{2}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}M_{H}^{2},&~{}~{}~{}T>T_{c},\\ \dfrac{1}{2}\Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}M_{H}^{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
and the W-boson mass which determines the monopole mass $M\simeq\bar{M}_{W}/\alpha$,
${\bar{M}}_{W}^{2}=\left\{\begin{array}[]{ll}0,&~{}~{}~{}T>T_{c},\\ \dfrac{g^{2}}{4}\rho_{+}^{2},&~{}~{}~{}T\leq T_{c}.\end{array}\right.$ (38)
The temperature dependent Higgs and W-boson masses are shown in Fig. 5.
Figure 5: The temperature dependent Higgs and W-boson masses. The blue and red curves represent the Higgs and W-boson masses.
Notice that the Higgs boson acquires the minimum mass $\bar{M}_{H}\simeq 5.5~{}{\rm GeV}$ at $T_{c}$ and approaches to the zero temperature mass 125.3 GeV as the universe cools down. Moreover, we have
$\bar{M}_{H}(T_{G})\simeq 116.9~{}{\rm GeV}.$ (39)
This tells that at $T_{G}$ already becomes close to the Higgs mass at zero temperature. But the W-boson is massless at high temperature above $T_{c}$ (before the symmetry breaking), and becomes massive toward the value 7.1 GeV at $T_{c}$ as the Higgs vacuum $\rho=0$ tunnels to $\rho_{+}(T_{c})\simeq 21.7~{}\text{GeV}$. Moreover, we have
$\bar{M}_{W}(T_{c})\simeq 7.1~{}{\rm GeV},~{}~{}~{}~{}\bar{M}_{W}(T_{G})\simeq 76.0~{}{\rm GeV}.$ (40)
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$), so that at $T_{G}$ the monopole mass acquires almost the adolecent value. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
The unique feature of the topological monopole production mechanism is that (unlike the Drell-Yan or Schwinger mechanism) the monopole mass depends on the temperature at creation, and that the baby monopole mass could be considerably smaller than the adolescent monopole mass. This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV.
This also confirms that what is important for the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition. The exponential suppression of the monopole production in the first order phase trabsition applies only when the Ginzburg temperature becomes higher than $T_{2}$. As far as the Ginzburg temperature becomes lower than $T_{2}$, there is no much difference between the monopole production in the first and second order phase transitions.
$T$ $\rho_{+}(T)$ $\bar{M}_{H}(T)$ $\bar{M}_{W}(T)$ $\bar{M}(T)$
$T_{1}$    $146.74$ $16.3$ $5.9$ $5.3$ $729.0$
$T_{c}$ $146.70$ $21.7$ $5.5$ $7.1$ $971.7$
$T_{2}$ $146.42$ $32.5$ $11.7$ $0.61$ 454.81454.81\,454.81 454.8
$T_{i}^{LHC}$ $132.89$ $119.2$ $56.8$ $38.9$ 5 331.55331.55\,331.55 331.5
$T_{min}^{LHC}$ $119.36$ $156.5$ $76.2$ $51.1$ 7 000.07000.07\,000.07 000.0
$T_{i}$ $101.96$ $188.4$ $92.9$ $61.5$ 8 428.28428.28\,428.28 428.2
$T_{G}$ $57.49$ $232.9$ $116.9$ $76.0$ 10 419.610419.610\,419.610 419.6
00 $0.00$       $246.2$       $125.3$       $80.4$ 11 014.511014.511\,014.511 014.5
Table 1: The values of $\rho_{+}$, $\bar{M}_{H}$, $\bar{M}_{W}$, and the expected monopole mass $\bar{M}=\bar{M}_{W}/\alpha$ at various temperatures. All numbers are in GeV.
IV Electroweak monopoles from a fireball at LHC
If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it is commonly believed, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does. In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
However, the 14 TeV LHC can not produce the monopole of mass 10.4 TeV. This means that the present LHC can not produce the monopole exactly as the early universe does. In other words, although LHC could reproduce the early universe, it could do so only in a limited sense. This is because, unlike the big bang fueled by the infinite energy, LHC can provide only a finite energy. So we have to take care of this energy constraint at LHC.
To do that, we estimate the temperature at which the monopole production stops at the present LHC. Since LHC should produce the monopole in pairs, the minimum temperature $T_{min}$ at the 14 TeV LHC for the monopole production must satisfy the energy condition
$M_{max}=7~{}\text{TeV}\simeq\frac{\bar{M}_{W}(T_{min})}{\alpha}=\frac{g}{2}\frac{\rho_{+}(T_{min})}{\alpha}.$ (41)
Solving this we find
$T_{min}^{LHC}\simeq 119.4~{}{\rm GeV}.$ (42)
This means that the energy condition to produce the monopole mass no more than 7 TeV forces the monopole production at LHC to stop at 119.4 GeV, much higher that the Ginzburg temperature. The effective potential at $T_{min}^{LHC}$ is also shown in Fig. 3 for comparison.
If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature. Moreover, in average the monopole production temperature at LHC is given by
$T_{i}^{LHC}=\frac{T_{c}+T_{min}^{LHC}}{2}\simeq 132.9~{}\text{GeV},$
$\rho_{+}(T_{i}^{LHC})\simeq 119.2~{}\text{GeV},$ (43)
not at $T_{i}$ given by (30).
This tells that LHC starts to produce the electroweak monopole with mass 1.0 TeV at around 146.7 GeV, and stops producing the monopole with mass 7 TeV at 119.4 TeV. In average LHC produces the infant electroweak monopole mass around 5.3 TeV, much less than the adolescent mass 11.0 TeV. This implies that the 14 TeV LHC could actually produce the electroweak monopole pair even when the mass of the monopole pair becomes bigger than 14 TeV. This is remarkable.
It is generally believed that in the Kibble-Zurek mechanism we are supposed to have one monopole per one correlation volume. This assumotion, however, may have a critical defect. This is because the correlation length is fixed by the electroweak scale but the monopole mass is given by the $1/\alpha$ times bigger than the electroweak scale, so that the energy in one correlation volume may not be enough to make up the monopole mass. A natural way to cure this defect is to enlarge the correlation length $\xi=1/\bar{M}_{H}$ to $\bar{\xi}$,
$\bar{\xi}=(\frac{1}{\alpha})^{1/3}~{}\xi\simeq 5.16\times\frac{1}{\bar{M}_{H}}.$ (44)
With this we have the new correlation volume at $T_{i}^{LHC}$,
$V_{c}\simeq\frac{4\pi^{2}}{3}\bar{\xi}^{3}\simeq 0.76\times 10^{-43}cm^{3}.$ (45)
In comparison the p-p fireball volume at the 14 TeV LHC is given by (with the proton radius $0.87\times 10^{-13}~{}cm$)
$V_{pp}\simeq\frac{0.938}{7000}\times\frac{4\pi^{2}}{3}~{}(0.87)^{3}\times 10^{-39}cm^{3}$
$\simeq 1.16\times 10^{-42}cm^{3},$ (46)
where the first term represents the Lorentz contraction of the p-p fireball. Notice that $V_{c}$ is about thirteen times bigger than $V_{pp}$. This tells that the p-p fireball at LHC has just enough size to produce the monopole.
We could translate the monopole production process at LHC in time scale. From (31) we can say that the electroweak monopole production at LHC starts at 146.7 GeV for the period of $0.7\times 10^{-11}~{}sec$. This tells that the monopole production lasts only very short period at LHC. So one might wonder if we have enough thermal fluctuations of the Higgs vacuum during this period. We can estimate how many times the Higgs vacuum fluctuates from $\rho_{+}(T_{LHC})$ to zero in average. From the uncertainty principle the time $\Delta t$ for one fluctuation is given by
$\Delta t\simeq\frac{1}{\Delta E}\simeq 4.7\times 10^{-27}~{}sec,$ (47)
so that the number of the fluctuation $N_{f}$ of the Higgs vacuum is given by
$N_{f}\simeq\frac{\bar{t}}{\Delta t}\simeq 3.1\times 10^{16}.$ (48)
This assures that we have enough fluctuations of the Higgs vacuum to produce the monopoles at LHC.
The above discussion tells that LHC could produce the electroweak baby monopole which could have the mass considerably less than the adolescent mass of 11.0 TeV, if we assume that the monopole production mechanism at LHC is topological. This is unexpected.
If so, can we really detect the infant electroweak monopole at LHC whose mass is smaller than the adolescent mass? To answer this we have to keep in mind that in the early universe most of the monopole-antimonopole pairs are quickly annhilated as soon as they are produced. This is because at the production temperature the capture radius of the monopole and anti-monopole is almost hundred times bigger than the correlation length. As the result most of the monopole-antimonopole pairs are annihilated as soon as they are produced, and only about $10^{-5}$ of the monopoles are survived in the early universe [27]. So one might wonder if we could have the same annihilation of the monopole-antimonopole pairs at LHC.
Actually, this annhilation is not expected to play an important role at LHC. This is because (unlike the early universe) the size of the p-p fireball at LHC is only about thirteen times bigger than the correlation volume, so that many of the monopoles could quickly escape the fireball before they are annhilated. This must be clear from (45) and (46). Moreover, once they escape from the fireball they are quickly freezed out (i.e., “quenched”). And as far as we know, nothing can destroy these quenched monopoles after they escape from the fireball. In this case MoEDAL should be able to detect these infant monopoles in principle.
However, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball. This necessiate us to study the monopolium bound state which we discuss in the following.
V Electroweak Monopolium: Atomic Model
To discuss the atomic model of the monopolium we let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
$\Big{(}-\frac{1}{2\mu}\nabla^{2}+V\Big{)}\Psi=E~{}\Psi,$
$V=-\frac{4\pi}{e^{2}}\frac{1}{r},$ (49)
where $\mu=M/2$ is the reduced mass and $V$ is the Coulombic magnetic potential of the monopole-antimonopole pair. Obviously this becomes formally identical to the Schrödinger equation of the Hydrogen atom, except that the coupling strenth of the potential is replaced by $4\pi/e$ and the electron mass is replaced by $\mu$.
From this we have the energy spectrum
$E_{n}=-\frac{M}{4\alpha^{2}n^{2}}=-\frac{4,692.3}{n^{2}}\times M$
$\simeq-\frac{50,207}{n^{2}}~{}{\rm TeV},$ (50)
where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 10.7~{}\text{TeV}$. So the Rydburg energy of the monopolium becomes 50,207 TeV. Moreover, the Bohr radius $R$ of the monopolium is given by
$R=\frac{\alpha}{\mu}\times\alpha^{2}=\frac{2\alpha^{3}}{M}\simeq\frac{5.6\times~{}10^{-9}}{M_{W}}$
$\simeq 1.4\times 10^{-24}~{}\mbox{cm},$ (51)
Notice, however, that these results are totally unreasonable. First, the above Rydburg energy is too big (roughly 2,346 times bigger than two monopole masses), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium. Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size.
This could mean the followings. First, the monopolium quickly annihilates itself, long before it reaches to the ground state. In fact (50) suggests that the monopolium could exist only at highly excited levels, probably with $n\geq 49$, because below this level the monopole-antimonopole mass can not provide the necessary binding energy. Or else, this could mean that the naive atomic description of the monopolium is not acceptable. In this case we have to find a more realistic model of the monopolium.
The reason for the above results originates from two facts. First, the magnetic coupling of the monopole is too strong. Second, the monopole mass is too big. So we could make the atomic model more realistic by improving on these points. For example, we could modify the magnetic Coulombic potential at short distances for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$. Moreover, we could make the monopole mass much less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
On the other hand, the above exercise has one positive side, because it implies that the binding energy of the electroweak monopolium could be big, of the order of TeV. This could make the monopolium mass considerably lighter than $2M$, so that we could have the monopolium of mass less than 14 TeV even when the mass of monopole-antimonopole pair is heavier than 14 TeV. With these modifications, we could have a more realistic model of the electroweak monopolium that we discuss now.
VI Electroweak Monopolium: A Realistic Model
To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, because the monopole (and antimonopole) has the size of $1/M_{W}$. This means that at short distances the singularity of the magnetic potential in (49) should be regularized.
So we modify the potential $V$ at short distance with a harmonic oscillator potential smoothly connected at $r_{0}$, and let
$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ \\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
where $r_{0}$ is of the order of $1/M_{W}$. The potential is shown in Fig. 6 in the unit of $r_{0}$.
Figure 6: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium is shown as a function of $r/r_{0}$.
The significance of this potential is that the coupling constant $k=2\pi/e^{2}r_{0}$ of the hamornic oscillator potential is automatically fixed by the smoothness of the potential at $r_{0}$. This immediately tells that the energy gap of the lowlying energy spectrum is given (with $r_{0}\simeq 1/M_{W}$) roughly by $M_{W}/2\alpha=M/2$, which reduces the energy gap of (50) by the factor $10^{-3}$. This implies that the modified potential (55) could describe a realistic electroweak monopolium. With this observation we could solve the Schrödinger equation keeping the reduced mass $\mu$ of the monopole-antimonopole pair arbitrary, allowing $\mu$ to have a small mass.
The first step is the usual separation of variables for wave functions in the regions $r<r_{0}$ (region I) and $r>r_{0}$ (region II) as
$\Psi_{\rm{I,II}}(r,\varphi,\theta)\equiv\frac{R_{\rm{I,II}}(r)}{r}Y_{l}^{m}(\varphi,\theta)\,,$ (56)
where $Y_{l}^{m}(\varphi,\theta)$ are the spherical harmonics. Next, we adopt dimensionless coordinates
$r=\frac{2\alpha}{M}x\,~{}~{}~{}~{}r_{0}=\frac{2\alpha}{M}x_{0},~{}~{}~{}~{}E\equiv-\frac{M}{4\alpha^{2}}\epsilon,$ (57)
where we denoted dimensionless binding energy as $\epsilon>0$. In these coordinates, the relevant Schrödinger equations for the radial parts read
$\ddot{R}_{\rm I}-\Big{(}\frac{l(l+1)}{x^{2}}+\frac{x^{2}}{x_{0}^{3}}-\frac{3}{x_{0}}+\epsilon\Big{)}R_{\rm I}=0,$
$\ddot{R}_{\rm II}-\Big{(}\frac{l(l+1)}{x^{2}}-\frac{2}{x_{0}}+\epsilon\Big{)}R_{\rm II}=0.$ (58)
Both of these can be mapped to the Kummer’s equation,
$z\frac{d^{2}w(z)}{dz^{2}}+(b-z)\frac{dw(z)}{dz}-aw(z)=0,$ (59)
by suitable choices of $z\equiv z(x)$ and parameters $a$ and $b$.
In particular, this mapping is facilitated via assignments,
$R_{\rm I}=x^{l+1}\exp\Big{(}-\frac{x^{2}}{2x_{0}^{3/2}}\Big{)}w(x^{2}/x_{0}^{3/2}),$
$a_{\rm I}=\frac{2l+3+x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4},~{}~{}~{}~{}b_{\rm I}=l+3/2,$
$R_{\rm II}=x^{l+1}\exp\big{(}-\sqrt{\epsilon}x\big{)}w(2\sqrt{\epsilon}x),$
$a_{\rm II}=l+1-\epsilon^{-1/2},\quad b_{\rm II}=2l+2.$ (60)
As is well-known, the Kummer’s equation has two independent solutions, the confluent hypergeometric function $M(a,b,z)\equiv\,_{1}F_{1}\bigl{(}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{|}z\bigr{)}$ and the Tricomi’s function $U(a,b,z)$.
In the region I ($x<x_{0}$), we need to satisfy the regularity condition at the origin $x=0$, which force us to exclude $U(a,b,z)$ as it has a singularity at $z=0$ in contrast to $M(a,b,z)$, which is an entire function in $z$. Hence, the general solution in the first region that is regular at the origin is given by
$R_{\rm I}=N_{\rm I}\,x^{l+1}\exp\Big{(}-\frac{x^{2}}{2x_{0}^{3/2}}\Big{)}$
×$\times_{1}F_{1}\Big{(}\begin{matrix}\frac{l}{2}+\frac{3}{4}+\frac{x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4}\\ l+\frac{3}{2}\end{matrix}\Bigr{|}\tfrac{x^{2}}{x_{0}^{3/2}}\Big{)},$ (63)
where $N_{\rm I}$ is an arbitrary constant at the moment.
In the region II ($x>x_{0}$), we need to impose boundary condition at $x\to\infty$ so that the resulting wave equation is square-integrable. This force us to discard $M(a,b,z)$ that behaves as $M(a,b,z)\sim\Gamma(b)\exp(z)~{}z^{a-b}/\Gamma(a)$ for large $z$ (in a certain wedge of the complex plane). Notice that this asymptotic behaviour holds only if $a$ is not a negative integer (and hence $\Gamma(a)$ blows up), which is exactly the case for the hydrogen atom. For our modified potential, however, the spectrum differs and hence $a\not=-k$, where $k$ is a positive integer.
On the other hand, we have
$U(a,b,x)\sim x^{-a}\,_{2}F_{0}\bigl{(}\begin{matrix}a&a-b+1\\ &\end{matrix}\bigr{|}-1/x\bigr{)},$ (66)
as $x$ goes to infinity. Hence, the general solution in the second region has a following form,
$R_{\rm II}=N_{\rm II}\,x^{l+1}\exp\big{(}-\sqrt{\epsilon}x\big{)}$
×$\times U\Big{(}\begin{matrix}l+1-\frac{1}{\sqrt{\epsilon}}\\ 2l+2\end{matrix}\Bigr{|}2\sqrt{\epsilon}x\Big{)},$ (69)
where $N_{\rm II}$ is again an arbitrary constant.
Figure 7: The radial monopolium wave function R(x) for $n=1,2,3,4,5$ shown in (63), (69), and (8).
The remaining task is to formulate sewing condition that ensures $C_{1}$ continuity of the wavefunction at the point $x=x_{0}$. The condition that $R_{\rm I}(x_{0})=R_{\rm II}(x_{0})$ fixes, say, $N_{\rm II}$ in terms of $N_{\rm I}$, which can be in turn fixed by square integrability condition. The other condition $R_{\rm I}^{\prime}(x_{0})=R_{\rm II}^{\prime}(x_{0})$ then boils down to a transcendental equation for the binding energy,
$\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{3+2l}\frac{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{7+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{5}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{3}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}$ (74)
$+2\sqrt{x_{0}}\big{(}(1+l)\sqrt{\epsilon}-1\big{)}\frac{U\Big{(}\begin{matrix}2+l-\epsilon^{-1/2}\\ 3+2l\end{matrix}\Big{|}2x_{0}\epsilon^{1/2}\Big{)}}{U\Big{(}\begin{matrix}1+l-\epsilon^{-1/2}\\ 2+2l\end{matrix}\Bigr{|}2x_{0}\epsilon^{1/2}\Big{)}}$ (79)
$+\sqrt{\epsilon x_{0}}=1.$ (80)
With the two conditions we have the smooth radial wave function of the monopolium. This is shown in Fig. 7.
Figure 8: The contours where Eq. (8) holds as function of $\sqrt{x_{0}}$ and $\sqrt{\epsilon}$ with $l=0$. The dashed lines represent harmonic oscillator levels given by (82). The horizontal dotted lines are the levels of the hydrogen atom, i.e., $\epsilon=1/n^{2}$, with $n=1,2,\ldots$
Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (8). Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig. 8 (for $l=0$).
We see that as $x_{0}$ goes to zero, the binding energies approaches the values given by the atomic model
$\epsilon_{A}\simeq\frac{1}{n^{2}},~{}~{}~{}~{}n=1,2,\ldots,$ (81)
whereas for large $x_{0}$, the contours quickly settles on the curves given by the binding energies of the harmonic oscillator, i.e.,
$\epsilon_{HO}=\frac{3\sqrt{x_{0}}-3-2l-4n}{x_{0}^{3/2}},~{}~{}~{}~{}n=0,1,\ldots$ (82)
The value of $x_{0}$ we are interested is given by $r_{0}\sim 1/M_{W}$, or
$x_{0}=\frac{M}{2\alpha}r_{0}\simeq 7,767,173.$ (83)
For such a high values of $x_{0}$, the binding energies are almost indistinguishable from binding energies of harmonic oscillator.
This means that the ground state energy of the monopolium (for $l=0$) is roughly
$E_{0}=-\frac{\epsilon_{0}}{4\alpha^{2}}~{}M\simeq-4,692\times(M/\text{TeV})~{}\epsilon_{HO}^{n,l=0}$
$\simeq-16.3~{}\text{TeV}.$ (84)
where we have put $M\simeq 10.7~{}\text{TeV}$. Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium. In particular, this tells that the monopole-antimonopole pair with mass 21.4 TeV makes the monopolium bound state of mass of 5.1 TeV at the ground state. This is precisely what we have hoped for.
Furthermore, the Bohr radius of the monopolium, which can be read of from the exponential decay of $R_{\rm II}$ from (69) is given by
$R=\sqrt{\frac{1}{|E_{0}|}}\simeq 1.5\times 10^{-18}~{}cm.$ (85)
This is $0.6\times 10^{-2}$ times smaller than the monopole size given by $1/M_{W}\simeq 2.5\times 10^{-16}~{}cm$. Again, this is definitely much more realistic than (51). This confirms that the monopolium bound state with the modified potential (6) can give us the desired binding energy which could help LHC to produce the bound state even when the monopole mass becomes bigger than 7 TeV.
It should be mentioned that a different model of the monopolium which assumes an exponential repulsion at short distance given by the potential
$V=-\frac{4\pi}{e^{2}}\frac{1}{r}\Big{(}1-\exp(-r/2r_{0})\Big{)},$ (86)
has been discussed before [33]. Despite the obvious difference between our potential (55) and this, we notice that the characteristic features of the two monopoliums are quite similar. In particular, in both cases the monopolium masses are slightly below the half of the monopole’s mass and the radiuses are of the order of $10^{-18}cm$. This implies that both could be used as a realistic model of the monopolium.
VII Discussion
An urgent task at LHC is to discover the electroweak monopole. Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it. But the problem is that the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV. And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, LHC cannot produce it. In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition.
Our work in this paper tells that, if LHC produces the monopole by the topological thermal fluctuation of the Higgs vacuum, LHC could circumbent this energy constraint and produce the electroweak monopole even when the monopole mass becomes 11.0 TeV. This is because in this thermal production the monopole mass depends on the temperature at production. In this case the mass of the monopoles produced right after the electroweak phase transition (the baby monopole mass) could be as low as 1.0 TeV, even when the monopole mass at zero temperature (the adolescent monopole mass) becomes 11.0 TeV.
So, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum (the same mechanism as the monopole production in the early universe), the present 14 TeV LHC could produce the electroweak monopole, and thus MoEDAL could detect it. Our results suggest that LHC could produce the infant monopoles of mass around 5.3 TeV in average (assuming the adolescent monopole mass is 11 TeV). This is really remarkable, because this implies that there is a finite possibility that the MoEDAL detector at LHC could actually detect the infant electroweak monopoles, even though it is not likely to detect the adolescent (i.e., grown-up) monopoles.
It should be emphasized that this is a unique feature of the finite temperature physics which is absent in the zero temperature physics. In particular, this is impossible if the monopole production mechanism at LHC is the Drell-Yan or Schwinger mechanism, where only the adolescent monopole enters. From this we can definitely say that the detection of the infant monopole at LHC should be interpreted as the confirmation that the electroweak monopole production mechanism is indeed topological.
Another important result of our work is that there is a real possibility for LHC to produce the monopoles in the form of the monopolia. When LHC produces the monopoles, it should produce them in pairs. But since they have strong magnetic attraction some of them could form bound states, although many of the monopole-antimonopole pairs could quickly annihilate. An interesting feature of this monopolium is that the mass could be considerably lighter than the mass of the monopole-antimonopole pair, because the monopolium binding energy could be of the order of ten TeV. In fact, our study in this paper suggests that the binding energy of the ground state could be as much as 16.7 TeV, so that the monopole-antimonopole pair of the mass 22.0 TeV could form a bound state of the mass 5.3 TeV. This provides another way for LHC to circumbent the energy constraint and produce the monopole pairs.
This tells that there are two ways for LHC to produce the monopole pairs, the monopole and antimonopole as independent particles or the monopole-antimonopole pair together as a monopolium bound state. So MoEDAL and ATLAS could try to detect the monopole (and antimonopole) separately, or together as a monopolium.
Detecting the monopole (and antimonopole) should be straightforward in principle, but detecting the monopolium could be tricky. This is because the monopolium has zero magnetic charge, so that we have to look for the magnetically neutral particles of mass of few TeV. Moreover, the monopole and antimonopole inside the monopolium could easily annihilate to photons and electron-positron pairs. In this case we may have to look for the decay modes of the monopolium.
One might wonder if we could also have the infant monopoles in cosmology produced in the early universe. This may not be possible, because the monopoles in the early universe are in contact with the hot thermal radiation bath which feeds up the energy and allows them to grow up. This means that the remnant monopoles from the early universe that we could detect at present are expected to be the adolescent monopoles. This implies that the quenched infant monopole at LHC is not a generic phenomenon but an outcome of the limited energy and size of the p-p fireball at LHC.
Admittedly the idea of the infant monopole at LHC sounds strange and appears hard to swallow. We emphasize, however, that this idea actually comes from the fact that in the electroweak phase transition the W (and Higg) boson mass becomes temperature dependent. Once we accept this, the idea of the infant monopole follows naturally. Moreover, it appears that (as far as we know) logically it is difficult to exclude this possibility. Ultimately, of course, only experiment could tell whether this idea is realistic or not.
We hope that our results in this paper could help MoEDAL to detect the electroweak monopole at LHC.
ACKNOWLEDGEMENT
PB and FB are supported by the Institute of Experimental and Applied Physics, Czech Technical University in Prague and the Research Centre for Theoretical Physics and Astrophysics, Institute of Physics, Silesian University in Opava. YMC is supported in part by the National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University, Korea. The authors thank their supports.

### v2

@@
 However, the present LHC cannot produce the monopole if the mass is bigger than 7 TeV, while the monopole mass is expected to be around $M_{W}/\alpha\simeq 11~{}\text{TeV}$.
-In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass is bigger than 7 TeV, based on the following ideas.
+In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass becomes bigger than 7 TeV, based on the following ideas.
 First, in the topological production of the monopole the baby monopole mass at creation could be considerably smaller than the adolescent mass.
@@
 We discuss how these ideas can actually be realized at LHC to produce the monopole pairs.
-In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.2 TeV, smaller than the adolescent monopole mass around 11.0 TeV.
-Moreover, we show that LHC could produce the monopolium bound state with mass around 5.3 TeV, even when the total mass of the monopole-antimonopole pair is around 22.0 TeV.
+In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.3 TeV, smaller than the adolescent monopole mass around 11.0 TeV.
+Moreover, we show that LHC could produce the monopolium bound state with mass around 2.5 TeV, even when the total mass of the baby monopole-antimonopole pair becomes around 10.6 TeV.
 Our analysis could play an important role for MoEDAL experiment.
@@
 However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV.
-This is problematic, because the monopole mass coud turn out to be bigger than this, around 11.0 TeV.
+This is problematic, because the monopole mass could turn out to be bigger than this, around 11.0 TeV.
 If this is the case, we may have little chance to detect the monopole at LHC.
@@
 This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole, or the quantized magnetic flux of the monopole-antimonopole pair.
-In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass is bigger than 7 TeV.
+In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
 The second point is related to the electroweak monopole production mechanism at LHC.
@@
 LHC could produce the monopole with this mechanism because the fireball of the p-p collision could re-create the electroweak phase transition in the early universe, and thus produces the monopoles topologically by the thermal fluctuation of the Higgs vacuum.
-A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass is heavier than 7 TeV.
+A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass becomes heavier than 7 TeV.
 To understand this point we first have to keep in mind that the monopole mass we have been talking about is the mass at the zero temperature.
@@
 This tells that we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass.
-So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
-The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV.
-We show that in the topological thermal production the monopole production at LHC starts from $T_{c}\simeq 147~{}\text{TeV}$ and stops at $T_{min}^{LHC}\simeq 119~{}\text{TeV}$ before the Ginzburg temperature $T_{G}\simeq 57.5~{}\text{GeV}$, with the monopole mass around 5.3 TeV.
-Moreover, we show that LHC could produce the monopolium bound states of mass around 5.3 TeV, even though the mass of the monopole-antimonopole pair becomes 22.0 TeV.
+So, from now on we will call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
+[…]
 As we will see, just after the phase transition the baby monopole mass could be as low as 1.0 TeV.
-Notice that this would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist.
+This would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist.
 On the other hand, we emphasize that the monopole production in the early universe and the monopole production at LHC have important differences.
@@
 This implies that LHC might need more energy to produce the magnetic flux made of the monopole-antimonopole pair than to produce the monopole and antimonopole separately.
-This makes the monopolium bound state another probable signal for the monopole production at LHC.
+This makes the monopolium bound state a more probable signal for the monopole production at LHC.
 The monopolium itself has been studied before for various reasons [32, 33, 34].
@@
 First, the naive atomic model of monopolium does not describe a realistic monopolium, because it has unrealistically big binding energy and too small size.
-This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is huge.
+This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is too big.
 But we could avoid these difficulties and have a realistic electroweak monopolium modifying the magnetic potential at short distance.
@@
 $=\left(\begin{array}[]{cc}\cos\theta_{\rm w}&\sin\theta_{\rm w}\\ -\sin\theta_{\rm w}&\cos\theta_{\rm w}\end{array}\right)\left(\begin{array}[]{cc}B_{\mu}\\ A_{\mu}^{\prime}\end{array}\right),$ (17)
-and have the the identity
+and have the identity
 $|{\cal D}_{\mu}\xi|^{2}=\frac{1}{4}Z_{\mu}^{2}+\frac{g^{2}}{2}W_{\mu}^{*}W_{\mu}.$ (18)
@@
 This is the gauge independent abelianization of the Weinberg-Salam Lagrangian.
-Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson masses are given by $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
+Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson acquire the mass $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
 The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field.
@@
 We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [23, 24, 25, 27, 28].
-In the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
+But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
 Figure 1: The Feynman diagrams of the popular monopole production mechanism at LHC given by Drell-Yan and two-photon fusion process. Here the proton pair could also be interpreted as the quark pair.
 To understand the electroweak monopole production at LHC we have to know the monopole production mechanism.
-There are two contending monopole production mechanisms in particle physics community at present, the Drell-Yan (and two photon fusion) process and the Schwinger mechanism.
 In the Drell-Yan process the monopole-antimonopole pairs are produced by the electromagnetic interaction through the photon [19, 20].
 But in the Schwinger mechanism a strong magnetic field creates the monopole-antimonopole pairs, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [22, 29, 30].
-The justification for this comes from the fundamental assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ with $(e,g)$ to $(g,-e)$.
+The justification for this comes from the assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$.
 The Drell-Yan process is shown graphically in Fig.
@@
 1.
-A unique feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant.
+A remarkable feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant.
 Perhaps more importantly, in this thermal fluctuation the monopole mass at creation (the baby monopole mass) could be considerably smaller than the adolescent monopole mass.
@@
 This large fluctuation disappears when the equality holds, at the Ginzburg temperature $T_{G}$.
-We can find it graphically from (21) and (28). This is shown in Fig. 4. From this we have [27]
+We can find $T_{G}$ graphically from (21) and (28). This is shown in Fig. 4. From this we have [27]
 $T_{G}\simeq 57.49~{}\text{GeV},~{}~{}~{}~{}\rho_{+}(T_{G})\simeq 232.9~{}\text{GeV}.$ (29)
@@
 3.
-Notice that at $\rho_{+}(T_{G})$ is quite close to the Higgs vacuum at zero temperature, which confirms that the electroweak monopole production lasts long time after the phase transition.
+Notice that $\rho_{+}(T_{G})$ is quite close to the Higgs vacuum at zero temperature, which confirms that the electroweak monopole production lasts long time after the phase transition.
 Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the black line represents $T_{G}$.
@@
 $\bar{M}_{H}(T_{G})\simeq 116.9~{}{\rm GeV}.$ (39)
-This tells that at $T_{G}$ already becomes close to the Higgs mass at zero temperature.
+This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature.
 But the W-boson is massless at high temperature above $T_{c}$ (before the symmetry breaking), and becomes massive toward the value 7.1 GeV at $T_{c}$ as the Higgs vacuum $\rho=0$ tunnels to $\rho_{+}(T_{c})\simeq 21.7~{}\text{GeV}$.
@@
 $\bar{M}_{W}(T_{c})\simeq 7.1~{}{\rm GeV},~{}~{}~{}~{}\bar{M}_{W}(T_{G})\simeq 76.0~{}{\rm GeV}.$ (40)
-This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$), so that at $T_{G}$ the monopole mass acquires almost the adolecent value.
+This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$).
+This tells that the baby monopole mass can be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value.
+This could allow LHC to produce the monopole even when the adolescent mass becomes bigger than 7 TeV.
 In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
-The unique feature of the topological monopole production mechanism is that (unlike the Drell-Yan or Schwinger mechanism) the monopole mass depends on the temperature at creation, and that the baby monopole mass could be considerably smaller than the adolescent monopole mass.
-This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV.
 This also confirms that what is important for the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
@@
 As far as the Ginzburg temperature becomes lower than $T_{2}$, there is no much difference between the monopole production in the first and second order phase transitions.
+At this point one might wonder if we could also observe the Higgs and W bosons which have the temperature dependent masses (the “baby” Higgs and W bosons) at LHC.
+This is not possible, because they are ordinary particles produced by the ordinary process, not the topological particles produced topologically.
+So there is no reason to believe that LHC can produce them.
 $T$ $\rho_{+}(T)$ $\bar{M}_{H}(T)$ $\bar{M}_{W}(T)$ $\bar{M}(T)$
@@
 To do that, we estimate the temperature at which the monopole production stops at the present LHC.
-Since LHC should produce the monopole in pairs, the minimum temperature $T_{min}$ at the 14 TeV LHC for the monopole production must satisfy the energy condition
-$M_{max}=7~{}\text{TeV}\simeq\frac{\bar{M}_{W}(T_{min})}{\alpha}=\frac{g}{2}\frac{\rho_{+}(T_{min})}{\alpha}.$ (41)
+Since LHC should produce the monopole in pairs, the minimum temperature $T_{min}^{LHC}$ at the 14 TeV LHC for the monopole production must satisfy the energy condition
+$M_{max}=7~{}\text{TeV}\simeq\frac{\bar{M}_{W}(T_{min}^{LHC})}{\alpha}=\frac{g}{2}\frac{\rho_{+}(T_{min}^{LHC})}{\alpha}.$ (41)
 Solving this we find
@@
 If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature.
-Moreover, in average the monopole production temperature at LHC is given by
+In average the monopole production temperature at LHC is given by
 $T_{i}^{LHC}=\frac{T_{c}+T_{min}^{LHC}}{2}\simeq 132.9~{}\text{GeV},$
@@
 In this case MoEDAL should be able to detect these infant monopoles in principle.
-However, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball.
+On the other hand, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball.
 This necessiate us to study the monopolium bound state which we discuss in the following.
 V Electroweak Monopolium: Atomic Model
-To discuss the atomic model of the monopolium we let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
+Let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
 $\Big{(}-\frac{1}{2\mu}\nabla^{2}+V\Big{)}\Psi=E~{}\Psi,$
@@
 where $\mu=M/2$ is the reduced mass and $V$ is the Coulombic magnetic potential of the monopole-antimonopole pair.
-Obviously this becomes formally identical to the Schrödinger equation of the Hydrogen atom, except that the coupling strenth of the potential is replaced by $4\pi/e$ and the electron mass is replaced by $\mu$.
+Obviously this is formally identical to the Schrödinger equation of the Hydrogen atom, except that the coupling strenth of the potential is replaced by $4\pi/e$ and the electron mass is replaced by $\mu$.
 From this we have the energy spectrum
 $E_{n}=-\frac{M}{4\alpha^{2}n^{2}}=-\frac{4,692.3}{n^{2}}\times M$
-$\simeq-\frac{50,207}{n^{2}}~{}{\rm TeV},$ (50)
+$\simeq-\frac{51,615.3}{n^{2}}~{}{\rm TeV},$ (50)
 where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 11.0~{}\text{TeV}$.
-So the Rydburg energy of the monopolium becomes 50,207 TeV.
+So the Rydburg energy of the monopolium becomes 51.615 TeV.
 Moreover, the Bohr radius $R$ of the monopolium is given by
@@
 Second, the monopole mass is too big.
-So we could make the atomic model more realistic by improving on these points.
+We could make the atomic model more realistic by improving on these points.
 For example, we could modify the magnetic Coulombic potential at short distances for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$.
-Moreover, we could make the monopole mass much less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
+Moreover, we could make the monopole mass less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
 On the other hand, the above exercise has one positive side, because it implies that the binding energy of the electroweak monopolium could be big, of the order of TeV.
@@
 Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig.
-8 (for $l=0$).
+8 for $l=0$.
 We see that as $x_{0}$ goes to zero, the binding energies approaches the values given by the atomic model
@@
 $E_{0}=-\frac{\epsilon_{0}}{4\alpha^{2}}~{}M\simeq-4,692\times(M/\text{TeV})~{}\epsilon_{HO}^{n,l=0}$
-$\simeq-16.3~{}\text{TeV}.$ (84)
+$\simeq-16.8~{}\text{TeV}.$ (84)
 where we have put $M\simeq 11.0~{}\text{TeV}$.
 Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium.
-In particular, this tells that the monopole-antimonopole pair with mass 21.4 TeV makes the monopolium bound state of mass of 5.1 TeV at the ground state.
+In particular, this tells that the monopole-antimonopole pair with mass 22.0 TeV makes the monopolium bound state of mass of 5.2 TeV at the ground state.
 This is precisely what we have hoped for.
@@
 This confirms that the monopolium bound state with the modified potential (6) can give us the desired binding energy which could help LHC to produce the bound state even when the monopole mass becomes bigger than 7 TeV.
+Notice that in the above analysis we have assumed that the monopole mass is 11 TeV.
+However, at LHC the monopole mass at $T_{min}^{LHC}$ becomes 5.3 TeV, so that the monopolium at LHC is expected to be the monopole-antimonopole bound state with mass 10.6 TeV.
+This will change the energy spectrum and the Bohr radius of the monopolium.
+For example, $E_{0}$ in (84) should change to $E_{0}\simeq-8.1~{}\text{TeV}$, which reduces the monopolium mass to 2.5 TeV at LHC.
 It should be mentioned that a different model of the monopolium which assumes an exponential repulsion at short distance given by the potential

### v4

@@
 Abstract
-One of the urgent issues in high energy physics is the experimental confirmation of the electroweak monopole predicted by the standard model, and currently MoEDAL at LHC is actively searching for the monopole.
-However, the present LHC cannot produce the monopole if the mass is bigger than 7 TeV, while the monopole mass is expected to be around $M_{W}/\alpha\simeq 11~{}\text{TeV}$.
-In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass becomes bigger than 7 TeV, based on the following ideas.
-First, in the topological production of the monopole the baby monopole mass at creation could be considerably smaller than the adolescent mass.
-Second, the binding energy of the monopole-antimonopole pair could effectively reduce the mass of the bound state.
-We discuss how these ideas can actually be realized at LHC to produce the monopole pairs.
-In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.3 TeV, smaller than the adolescent monopole mass around 11.0 TeV.
-Moreover, we show that LHC could produce the monopolium bound state with mass around 2.5 TeV, even when the total mass of the baby monopole-antimonopole pair becomes around 10.6 TeV.
-Our analysis could play an important role for MoEDAL experiment.
-electroweak monopole, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
+The monopole production at LHC crucially depends on the monopole production mechanism.
+We show that, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum as the early universe did, it is practically impossible for LHC to produce the monopole, even with the future FCC energy upgrade.
+This is because the temperature of the p-p fireball is simply too low to produce the monopole.
+But if the monopole production mechanism is the Drell-Yan and/or Schwinger mechanism, the 14 TeV LHC could produce the monopole in the form of the monopolium of mass around 5,7 TeV even when the monopole mass becomes 11 TeV.
+Or, it could produce the monopole when the mass is less than 7 TeV.
+Our result tells that, if the monopole production mechanism at LHC becomes the thermal fluctuation, searching for the remnant monopoles produced in the early universe could be the only way to detect the electroweak monopole.
+We discuss the physical implications of our result.
+electroweak monopole, monopole production mechanism at LHC, Drell-Yan process, Schwiner mechanism, thermal fluctuation of Higgs vacuum, Ginzburg temperature, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
 I Introduction
-With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2, 3].
-After the Dirac monopole we have had the Wu-Yang monopole, the ’t Hooft-Polyakov monopole, the grand unification monopole [4, 5, 6, 7, 8, 9, 10, 11].
-But the electroweak monopole (also known as the Cho-Maison monopole), despite the initial skepticism on the existence of such monopole in the standard model [12, 13], stands out as the most realistic monopole that could exist in nature and could actually be detected [14, 15, 16].
+With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2].
+After the Dirac monopole we have had the Wu-Yang monopole, the ’t Hooft-Polyakov monopole, and the grand unification monopole [3, 4, 5].
+But the electroweak (Cho-Maison) monopole, stands out as the most realistic monopole that could exist in nature and could actually be detected [6, 7].
 Indeed the Dirac monopole in electrodynamics should transform to the electroweak monopole after the unification of the electromagnetic and weak interactions, and the Wu-Yang monopole in QCD is supposed to make the monopole condensation to confine the color.
-Moreover, the ’tHooft-Polyakov monopole exists only in an hypothetical theory, and the grand unification monopole which could have been amply produced at the grand unification scale in the early universe probably has become completely irrelevant at present universe after the inflation [17, 18].
+Moreover, the ’tHooft-Polyakov monopole exists only in an hypothetical theory, and the grand unification monopole which could have been amply produced at the grand unification scale in the early universe probably has become completely irrelevant at present universe after the inflation.
 This makes the experimental confirmation of the electroweak monopole one of the most urgent issues in the standard model after the discovery of the Higgs particle.
@@
 For this reason the newest MoEDAL detector at LHC is actively searching for the monopole [8, 9, 10, 11].
-To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [14, 15, 23, 24, 25, 26, 27, 28].
+To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [6, 7, 12, 13, 14, 15, 16].
 First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’t Hooft-Polyakov monopoles.
 Second, the magnetic charge of the monopole is not $2\pi/e$ but $4\pi/e$, twice that of the Dirac monopole.
-This is because the period of the electromagnetic U(1) subgroup of the standard model is $4\pi$, since the electromagnetic U(1) (partly) comes from the U(1) subgroup of SU(2).
-Third, the mass of the monopole $M$ is of the order of several TeV, probably between 4 to 10 TeV.
-This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the coupling is magnetic (i.e., $4\pi/e$).
-This makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV.
-Despite this, the size of the monopole is set by the weak boson masses.
-This is because the monopole solution has the Higgs and W boson dressing which has the exponential damping fixed by the weak boson masses.
+This is because the period of the electromagnetic U(1) subgroup of the standard model becomes $4\pi$.
+Third, the mass of the monopole $M$ is of the order of 10 TeV.
+This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the magnetic coupling makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV [17].
+Despite this, the size of the monopole is set by the W boson mass, because the monopole solution has the Higgs and W boson dressing which fixes the size by the W boson mass.
 Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it.
 However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV.
-This is problematic, because the monopole mass could turn out to be bigger than this, around 11.0 TeV.
-If this is the case, we may have little chance to detect the monopole at LHC.
-In this case we may have to try to detect the remnant monopoles at present universe produced in the early universe [27], or wait for the next upgrade of LHC.
-However, two observations could make the monopole signal at LHC non-zero even when the mass becomes over 7 TeV.
+This is problematic, because the monopole mass could turn out to be bigger than this.
+If this is so, we may have little chance to detect the monopole at LHC, and may have to try to detect the remnant monopoles at present universe produced in the early universe [15, 18].
+However, two observations could make the monopole production at LHC possible even when the mass becomes over 7 TeV.
 First, the monopole has to be created in pairs at LHC.
-This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole, or the quantized magnetic flux of the monopole-antimonopole pair.
+This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole.
 In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
 Two popular ones in the high energy physics community are the Drell-Yan (and two photon fusion) process [8, 9] and the Schwinger mechanism [10, 19, 20].
-The third one is the topological monopole production mechanism, in which the thermal fluctuation of the Higgs vacuum which takes place during the electroweak phase transition produces the monopole [27].
-LHC could produce the monopole with this mechanism because the fireball of the p-p collision could re-create the electroweak phase transition in the early universe, and thus produces the monopoles topologically by the thermal fluctuation of the Higgs vacuum.
-A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass becomes heavier than 7 TeV.
-To understand this point we first have to keep in mind that the monopole mass we have been talking about is the mass at the zero temperature.
-However, in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [27].
-This tells that we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass.
-So, from now on we will call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
-[…]
-As we will see, just after the phase transition the baby monopole mass could be as low as 1.0 TeV.
-This would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist.
-On the other hand, we emphasize that the monopole production in the early universe and the monopole production at LHC have important differences.
-First, the monopole production in the early universe stops around the Ginzburg temperature, while the monopole production at LHC stops at higher temperature.
-This is because at the present LHC the monopole mass can not exceed 7 TeV.
-So LHC has less monopole production time.
-Second, the size of the thermal radiation bath in the early universe is limited only by the horizon distance, but the size of the fireball at LHC in p-p collision is fixed by the proton size which is very small.
-Because of this, the monopole in the early universe has unlimited energy nearby but the monopole at LHC has very limited energy coming from the p-p fireball.
-These differences play important roles in the monopole production at LHC.
-The existence of a monopole-antimonopole bound state, the metastable quantized magnetic flux made of Cho-Maison monopole-antimonopole pair, has recently been confirmed [31].
-It has the expected pole separation length around $8.2/M_{W}$ and has the magnetic dipole moment roughly $8.6/M_{W}$.
-This extended size (length) of the magnetic flux suggests that the energy of the pair might be bigger than the sum of the monopole and antimonopole masses.
-This implies that LHC might need more energy to produce the magnetic flux made of the monopole-antimonopole pair than to produce the monopole and antimonopole separately.
-This makes the monopolium bound state a more probable signal for the monopole production at LHC.
-The monopolium itself has been studied before for various reasons [32, 33, 34].
-For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV, the mass scale of the low lying hadrons [34].
-But the electroweak monopolium made of the Cho-Maison monopole pair has not been studied before.
-Our study on the electroweak monopolium tells the followings.
-First, the naive atomic model of monopolium does not describe a realistic monopolium, because it has unrealistically big binding energy and too small size.
-This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is too big.
-But we could avoid these difficulties and have a realistic electroweak monopolium modifying the magnetic potential at short distance.
-As we will see, in this modification we can show that the binding energy reduces the monopolium mass from 22.0 TeV to 5.3 TeV.
-This is precisely what we hoped for.
-With this we could conclude that LHC could produce the electroweak monopolium of the mass considerably lighter than 14 TeV without much difficulty.
-In this case we could try to detect the monopolium bound state experimentally detecting the decay modes.
+The third one is the topological monopole production mechanism, in which the thermal fluctuation of the Higgs vacuum produces the monopole as the early universe did [15, 18].
+LHC could produce the monopole with this mechanism because the fireball of the p-p collision could reproduce the early universe.
+A unique feature of this topological thermal fluctuation mechanism is that in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [15, 18].
+In this case there is a possibility that LHC could produce the electroweak monopole even when the zero temperature mass becomes heavier than 7 TeV.
+The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV.
+We show that, if the monopole production mechanism at LHC is the Drell-Yan process (and/or Schwinger mechanism), the present 14 TeV LHC could produce the monopole if the mass is less than 7 TeV.
+Or, it could produce the monopole in the form of the monopolium bound states of mass around 5.7 TeV even when the monopole mass becomes 11 TeV.
+This is because the binding energy of the monopoliun could reduce the monopolium mass by 16.3 TeV.
+Indeed we argue that the LHC can produce the monopolium for any reasonable monopole mass, as far as the mass does not exceed 29.8 TeV.
+However, if the monopole production mechanism at LHC becomes the thermal fluctuation of the Higgs vacuum, it has practically no chance to produce the monopole even with the future FCC energy upgrade.
+This is because the temperature of the p-p fireball at LHC simply becomes too low to produce the monopole thermally.
+This makes the monopolium bound state a most probable signal for the monopole production at LHC.
 The paper is organized as follows.
@@
 In Section III we review the monopole production in the electroweak phase transition in the early universe.
-In Section IV we discuss how LHC could produce the monopole in a similar manner, and argue that LHC could produce the baby electroweak monopole whose mass is smaller than the adolescent mass.
-In Section V we discuss the naïve Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium.
+In Section IV we discuss if LHC could produce the monopole in a similar manner, and argue that this is practically impossible because the temperature of the p-p fireball is too low.
+In Section V we discuss the naive Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium.
 In Section VI we discuss a more realistic electroweak monopolium that could be produced at LHC.
-In Section VII we discuss the physical implications of our results.
-II Electroweak Monopole Production Mechanisms at LHC
-Before we discuss the electroweak monopole production at LHC, we review the standard model. Let us start from the (bosonic sector of the) Weinberg-Salam Lagrangian,
+In Section VII we discuss the physical impications of our results.
+II Electroweak Monopole Production Mechanism at LHC
+Consider the (bosonic sector of the) Weinberg-Salam Lagrangian,
 ${\cal L}_{WS}=-|{\cal D}_{\mu}\phi|^{2}-\frac{\lambda}{2}\big{(}\phi^{\dagger}\phi-\frac{\mu^{2}}{\lambda}\big{)}^{2}-\frac{1}{4}\vec{F}_{\mu\nu}^{2}-\frac{1}{4}G_{\mu\nu}^{2},$
 ${\cal D}_{\mu}\phi=\big{(}\partial_{\mu}-i\frac{g}{2}\vec{\tau}\cdot{\vec{A}}_{\mu}-i\frac{g^{\prime}}{2}B_{\mu}\big{)}\phi,$ (1)
-where $\phi$ is the Higgs doublet, $\vec{F}_{\mu\nu}$ and $G_{\mu\nu}$ are the gauge field strengths of $SU(2)$ and $U(1)_{Y}$ with the potentials ${\vec{A}}_{\mu}$ and $B_{\mu}$, and $g$ and $g^{\prime}$ are the corresponding coupling constants, and $\cal D_{\mu}$ is the covariant derivative of the $SU(2)$ and $U(1)_{Y}$ gauge groups.
+where $\phi$ is the Higgs doublet, $\vec{F}_{\mu\nu}$ and $G_{\mu\nu}$ with potentials ${\vec{A}}_{\mu}$ and $B_{\mu}$ are the gauge fields of $SU(2)$ and $U(1)_{Y}$, $\cal D_{\mu}$ is the covariant derivative, and $g$ and $g^{\prime}$ are the corresponding coupling constants.
 Expressing $\phi$ with the Higgs field $\rho$ and unit doublet $\xi$ by
@@
 $-g^{2}(W_{\mu}^{*}W_{\nu}-W_{\nu}^{*}W_{\mu})^{2},$
-$F^{\prime}_{\mu\nu}=\partial_{\mu}A^{\prime}_{\nu}-\partial_{\nu}A^{\prime}_{\mu},$
-$A^{\prime}_{\mu}=A_{\mu}+C_{\mu},~{}~{}~{}C_{\mu}=-\frac{2i}{g}\xi^{\dagger}\partial_{\mu}\xi$
+$F^{\prime}_{\mu\nu}=\partial_{\mu}A^{\prime}_{\nu}-\partial_{\nu}A^{\prime}_{\mu},~{}~{}A^{\prime}_{\mu}=A_{\mu}+C_{\mu},~{}~{}C_{\mu}=-\frac{2i}{g}\xi^{\dagger}\partial_{\mu}\xi$
 $D^{\prime}_{\mu}=\partial_{\mu}+igA^{\prime}_{\mu},~{}~{}~{}W_{\mu}=\frac{W^{1}_{\mu}+iW^{2}_{\mu}}{\sqrt{2}}.$ (5)
@@
 $|{\cal D}_{\mu}\xi|^{2}=\frac{1}{4}Z_{\mu}^{2}+\frac{g^{2}}{2}W_{\mu}^{*}W_{\mu}.$ (18)
+Notice that here we have defined $A_{\mu}^{(em)}$ and $Z_{\mu}$ gauge independently, without choosing the unitary gauge.
+Figure 1: The Feynman diagrams of the popular monopole production mechanism at LHC given by Drell-Yan and two-photon fusion process. Here the proton pair could also be interpreted as the quark pair.
 From this we can express the Weinberg-Salam Lagrangian in terms of the physical fields,
@@
 This is the gauge independent abelianization of the Weinberg-Salam Lagrangian.
-Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson acquire the mass $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
+Notice that here the Higgs doublet disappears completely, and the Higgs, W, and Z bosons acquire the mass $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$, without any spontaneous symmetry breaking.
+Figure 2: The topological monopole production mechanism in the early universe or at LHC induced by the thermal fluctuation of the Higgs vacuum.
 The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field.
-It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [14].
+It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [6, 7].
 Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV.
-Intuitively we could argue the mass to be $1/\alpha$ times heavier than the W boson mass, around 11 TeV.
+Intuitively we could argue the mass to be $1/\alpha$ times bigger than the W boson mass, around 11 TeV [17].
 This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$).
@@
 But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
-Figure 1: The Feynman diagrams of the popular monopole production mechanism at LHC given by Drell-Yan and two-photon fusion process. Here the proton pair could also be interpreted as the quark pair.
-To understand the electroweak monopole production at LHC we have to know the monopole production mechanism.
-In the Drell-Yan process the monopole-antimonopole pairs are produced by the electromagnetic interaction through the photon [19, 20].
-But in the Schwinger mechanism a strong magnetic field creates the monopole-antimonopole pairs, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [22, 29, 30].
-The justification for this comes from the assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$.
+To understand the electroweak monopole production at LHC we have to know the monopole production mechanism at LHC.
+In the Drell-Yan process the monopole-antimonopole are produced in pairs by the electromagnetic interaction through the photon [8, 9].
+But in the Schwinger mechanism a strong magnetic field is supposed to create the monopole-antimonopole pair, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [10, 19].
+The justification for this comes from the asumption that a theory of monopole must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$.
 The Drell-Yan process is shown graphically in Fig.
 1.
-Notice, however, that something is strange with these pictures.
-For instance, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process.
-But the Weinberg-Salam Lagrangian has no interaction which can be described by the above Feynman diagram.
+But in the topological monopole production, the thermal fluctuation of the Higgs vacuum induces the change of topology and produce the monopole, which is exactly the electroweak monopole production mechanism in the early universe [15].
+In this picture the monopoles are produced after the phase transition and stops at the Ginzburg temperature [15].
+The topological monopole production mecanism is shown graphically in Fig.
+2 for comparison.
+Notice that, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process.
+But the Weinberg-Salam Lagrangian perturbatively has no interaction which can be described by the above Feynman diagram.
 Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background.
-Most importantly, the monopole is not an ordinary point particle but a topological particle, so that we need a change of topology to produce it.
-But it is not clear how the Drell-Yan process or the Schwinger mechanism could accommodate this topological nature of the monopole.
-Figure 2: The topological monopole production mechanism in the early universe or at LHC induced by the thermal fluctuation of the Higgs vacuum.
-The standard model provides another monopole production mechanism, the topological monopole production generated by the thermal fluctuation of the Higgs vacuum which induces the change of topology, which is the electroweak monopole production mechanism in the early universe [27].
-In this picture the monopoles are produced in the course of the electroweak phase transition by thermal fluctuations of the Higgs vacuum, which takes place after the phase transition and stops at the Ginzburg temperature [27].
-This is shown in Fig.
-2, which should be contrasted with Fig.
-1.
-A remarkable feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant.
-Perhaps more importantly, in this thermal fluctuation the monopole mass at creation (the baby monopole mass) could be considerably smaller than the adolescent monopole mass.
-This is because in this thermal production, the monopole mass depends on the temperature at creation.
-As we will see in the following this could have a deep impact on the monopole production at LHC.
+Most importantly, we need the change of topology to produce the monopole, but this topological nature of the monopole has no role at all in these mechanisms.
+In comparison the topology plays an essential role in the thermal production of the monopole, because here the thermal fluctuation of the Higgs vacuum is precisely what we need to induce the change of topology.
+As a result the monopole production in this mechanism is not controlled by any fundamental interaction or fundamental constant.
+Perhaps more importantly, in this thermal fluctuation the monopole mass depends on the temperature at creation, the monopole mass at creation could be considerably smaller than the monopole mass at zero temperature.
+In this case we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass.
+So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” (or “infant”) monopole mass and the zero temperature mass as the “adolescent” monopole mass.
+It should be emphasized that the above monopole production mechanisms at LHC are the logical possibilities, not the realistic possibilities.
+For example, the Schwinger mechanism is certainly a logical possibility.
+However, it is not clear at all if (and how) the p-p fireball at LHC could create a strong magnetic background which can actually produce the monopole-antimonopole pair.
 If so, what is the the monopole production mechanism at LHC?
 We do not know yet.
-On the other hand, it has often been claimed that LHC should reproduce the early universe.
-If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production the early universe.
-In this paper we will adopt this view and discuss how LHC could produce the electroweak monopole.
-For this we have to understand the electroweak monopole production in the early universe.
+On the other hand, it has often been claimed that LHC could reproduce the early universe.
+If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production in the early universe.
+In the following we discuss if LHC could produce the electroweak monopole thermally.
 III Monopole Production in Early Universe: A Review
@@
 This tells that what is important in the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
-To amplify this point we start from the temperature dependent effective action of the standard model expressed by [27, 38, 39, 40, 41, 42, 43]
+To amplify this point we start from the temperature dependent effective action of the standard model which describes the electroweak monopole production in the early universe [15, 25, 26, 18]
 $V_{eff}(\rho)=V_{0}(\rho)-\frac{C_{1}}{12\pi}\rho^{3}~{}T+\frac{C_{2}}{2}\rho^{2}~{}T^{2}-\frac{\pi^{2}}{90}N_{*}T^{4}$
@@
 $C_{2}=\frac{4M_{W}^{2}+2M_{Z}^{2}+M_{H}^{2}+4m_{t}^{2}}{8\rho_{0}^{2}}\simeq 0.37,$ (21)
-where $V_{0}$ (with $\lambda\simeq 0.26$ and $\rho_{0}\simeq 246~{}{\rm GeV}$) is the zero-temperature potential, $N_{*}$ is the total number of distinct helicity states of the particles with mass smaller than $T$ (counting fermions with the factor 7/8), $C_{1}$ and $C_{2}$ are the contributions from the $\rho$ dependent part of the masses of gauge bosons, Higgs field, and fermions, $m_{t}$ is the top quark mass, and $\delta V_{T}$ is the slow-varying logarithmic corrections and the lighter quark contributions which we will neglect from now on.
+where $V_{0}$ (with $\lambda\simeq 0.26$ and $\rho_{0}\simeq 246~{}{\rm GeV}$) is the zero-temperature potential, $C_{1}$ and $C_{2}$ terms are the loop contributions from the gauge bosons, Higgs field, and heavy fermions, $N_{*}$ is the total number of distinct helicity states of the particles with mass smaller than $T$ (counting fermions with the factor 7/8), $m_{t}$ is the top quark mass, and $\delta V_{T}$ is the slow-varying logarithmic corrections and the lighter quark contributions which we will neglect from now on.
 The potential has three local extrema at
@@
 So above this temperature only $\rho_{s}=0$ becomes the true vacuum of the effective potential, and the electroweak symmetry remains unbroken.
+Figure 3: The effective potential (21) at different temperatures.
+Notice that the potential at $T_{1},~{}T_{c},~{}T_{2}$ are almost indistinguishable.
+Here the unit of $V_{eff}$ is chosen to be $V_{0}=(\lambda/8)\rho_{0}^{4}=1$.
 At $T=T_{1}$ we have
@@
 At $T_{c}$ the new vacuum bubbles start to nucleate at $\rho_{s}=0$, which tunnels to the stable vacuum $\rho_{+}$ after $T_{c}$.
-Below this critical temperature $\rho_{+}$ becomes the true minimum of the effective potential, but $\rho_{s}=0$ remains a local (unstable) minimum till the temperature reaches $T_{2}$,
+Below this critical temperature $\rho_{+}$ becomes the true minimum of the effective potential, but $\rho_{s}=0$ remains a local minimum till the temperature reaches $T_{2}$,
 $T_{2}=\sqrt{\frac{\lambda}{2C_{2}}}~{}\rho_{0}\simeq 146.42~{}\text{GeV},$
@@
 So for all practical purposes we could treat the electroweak phase transition as a second order phase transition.
-Figure 3: The effective potential (21) at different temperatures.
-Notice that the potentials at $T_{1},~{}T_{c},~{}T_{2}$ are almost indistinguishable.
-Here the unit of $V_{eff}$ is chosen to be $V_{0}=(\lambda/8)\rho_{0}^{4}=1$.
+Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the black line represents $T_{G}$.
 The monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism, so that the monopole production start from $T_{c}$.
-However, we have to keep in mind that the thermal fluctuations of the Higgs vacuum which create the seed of the monopoles continue as long as [27, 37]
+And the thermal fluctuations of the Higgs vacuum which create the seed of the monopoles continue as long as we have [15, 24]
 $\xi^{3}\Delta F\leq T,~{}~{}~{}~{}\Delta F(T)=V(\rho_{s})-V(\rho_{+}),$ (28)
@@
 Notice that $\rho_{+}(T_{G})$ is quite close to the Higgs vacuum at zero temperature, which confirms that the electroweak monopole production lasts long time after the phase transition.
-Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the black line represents $T_{G}$.
 With this observation we can say that the monopole formation takes place between $T_{c}$ and $T_{G}$, or roughly around $T_{i}$,
@@
 $\rho_{+}(T_{i})\simeq 188.4~{}\text{GeV}.$ (30)
-To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [44]
+To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [kolb]
 $t=\Big{(}\frac{90}{32\pi GN_{*}(T)}\Big{)}^{1/2}~{}\frac{1}{T^{2}}.$ (31)
@@
 From this we can say that the electroweak monopole production start from $1.8\times 10^{-11}sec$ to $1.2\times 10^{-10}sec$ after the big bang for the period of $10.3\times 10^{-11}~{}sec$, or around $3.5\times 10^{-11}sec$ after the big bang in average.
+Figure 5: The temperature dependent Higgs and W boson masses. The blue and red curves represent the Higgs and W boson masses.
 The effective potential (21) gives us two important parameters of the electroweak phase transition, the temperature dependent Higgs mass $\bar{M}_{H}$ which determine the correlation length $\xi=1/\bar{M}_{H}$,
 $\bar{M}_{H}^{2}=\frac{d^{2}V_{eff}}{d\rho^{2}}\Big{|}_{\rho_{min}}$
-=$=\left\{\begin{array}[]{ll}\dfrac{1}{2}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}M_{H}^{2},&~{}~{}~{}T>T_{c},\\ \dfrac{1}{2}\Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}M_{H}^{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
+=$=\left\{\begin{array}[]{ll}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T>T_{c},\\ \Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
 and the W-boson mass which determines the monopole mass $M\simeq\bar{M}_{W}/\alpha$,
@@
 The temperature dependent Higgs and W-boson masses are shown in Fig. 5.
-Figure 5: The temperature dependent Higgs and W-boson masses. The blue and red curves represent the Higgs and W-boson masses.
-Notice that the Higgs boson acquires the minimum mass $\bar{M}_{H}\simeq 5.5~{}{\rm GeV}$ at $T_{c}$ and approaches to the zero temperature mass 125.3 GeV as the universe cools down. Moreover, we have
-$\bar{M}_{H}(T_{G})\simeq 116.9~{}{\rm GeV}.$ (39)
-This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature.
-But the W-boson is massless at high temperature above $T_{c}$ (before the symmetry breaking), and becomes massive toward the value 7.1 GeV at $T_{c}$ as the Higgs vacuum $\rho=0$ tunnels to $\rho_{+}(T_{c})\simeq 21.7~{}\text{GeV}$.
-Moreover, we have
-$\bar{M}_{W}(T_{c})\simeq 7.1~{}{\rm GeV},~{}~{}~{}~{}\bar{M}_{W}(T_{G})\simeq 76.0~{}{\rm GeV}.$ (40)
-This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$).
-This tells that the baby monopole mass can be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value.
-This could allow LHC to produce the monopole even when the adolescent mass becomes bigger than 7 TeV.
-In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
-This also confirms that what is important for the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
-The exponential suppression of the monopole production in the first order phase trabsition applies only when the Ginzburg temperature becomes higher than $T_{2}$.
-As far as the Ginzburg temperature becomes lower than $T_{2}$, there is no much difference between the monopole production in the first and second order phase transitions.
-At this point one might wonder if we could also observe the Higgs and W bosons which have the temperature dependent masses (the “baby” Higgs and W bosons) at LHC.
-This is not possible, because they are ordinary particles produced by the ordinary process, not the topological particles produced topologically.
-So there is no reason to believe that LHC can produce them.
 $T$ $\rho_{+}(T)$ $\bar{M}_{H}(T)$ $\bar{M}_{W}(T)$ $\bar{M}(T)$
-$T_{1}$    $146.74$ $16.3$ $5.9$ $5.3$ $729.0$
+$T_{1}$    $146.74$ $16.3$ $5.9$ 00 00
 $T_{c}$ $146.70$ $21.7$ $5.5$ $7.1$ $971.7$
@@
 Table 1: The values of $\rho_{+}$, $\bar{M}_{H}$, $\bar{M}_{W}$, and the expected monopole mass $\bar{M}=\bar{M}_{W}/\alpha$ at various temperatures. All numbers are in GeV.
-IV Electroweak monopoles from a fireball at LHC
-If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it is commonly believed, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does.
+Notice that the Higgs boson acquires the minimum mass $\bar{M}_{H}\simeq 5.5~{}{\rm GeV}$ at $T_{c}$ and approaches to the zero temperature mass 125.3 GeV as the universe cools down. Moreover, we have
+$\bar{M}_{H}(T_{G})\simeq 116.9~{}{\rm GeV}.$ (39)
+This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature.
+But the W-boson which is massless above $T_{c}$ (before the symmetry breaking) becomes massive at $T_{c}$, and we have
+$\bar{M}_{W}(T_{c})\simeq 7.1~{}{\rm GeV},~{}~{}~{}~{}\bar{M}_{W}(T_{G})\simeq 76.0~{}{\rm GeV}.$ (40)
+This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$).
+This confirms that the baby monopole mass near $T_{c}$ could be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value.
+This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV.
+In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
+This also confirms that what is important for the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
+The exponential suppression of the monopole production in the first order phase trabsition applies only when the Ginzburg temperature becomes higher than $T_{2}$.
+As far as the Ginzburg temperature becomes lower than $T_{2}$, there is no much difference between the monopole production in the first and second order phase transitions.
+IV Electroweak Monopole Production from p-p Fireball at LHC
+If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it has been asserted, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does.
 In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
@@
 This assures that we have enough fluctuations of the Higgs vacuum to produce the monopoles at LHC.
-The above discussion tells that LHC could produce the electroweak baby monopole which could have the mass considerably less than the adolescent mass of 11.0 TeV, if we assume that the monopole production mechanism at LHC is topological.
-This is unexpected.
-If so, can we really detect the infant electroweak monopole at LHC whose mass is smaller than the adolescent mass?
-To answer this we have to keep in mind that in the early universe most of the monopole-antimonopole pairs are quickly annhilated as soon as they are produced.
-This is because at the production temperature the capture radius of the monopole and anti-monopole is almost hundred times bigger than the correlation length.
-As the result most of the monopole-antimonopole pairs are annihilated as soon as they are produced, and only about $10^{-5}$ of the monopoles are survived in the early universe [27].
-So one might wonder if we could have the same annihilation of the monopole-antimonopole pairs at LHC.
-Actually, this annhilation is not expected to play an important role at LHC.
-This is because (unlike the early universe) the size of the p-p fireball at LHC is only about thirteen times bigger than the correlation volume, so that many of the monopoles could quickly escape the fireball before they are annhilated.
-This must be clear from (45) and (46).
-Moreover, once they escape from the fireball they are quickly freezed out (i.e., “quenched”).
-And as far as we know, nothing can destroy these quenched monopoles after they escape from the fireball.
-In this case MoEDAL should be able to detect these infant monopoles in principle.
-On the other hand, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball.
-This necessiate us to study the monopolium bound state which we discuss in the following.
+The above discussion tells that LHC could produce the electroweak baby monopole of mass around 5.3 TeV even when the adolescent mass becomes 11.0 TeV, if we assume that the monopole production mechanism at LHC is thermal.
+If this is true, this would be really remarkable.
+Unfortunately, the above argument has a critical defect, because here we have implicitly assumed that the temperature of the p-p fireball would be high enough to produce the monopole thermally (assuming that the 14 TeV p-p fireball energy would produce the high temperature necessary for LHC to produce the monopole thermally).
+But this is not obvious, because the temperature and energy of the fireball are two different things.
+As we have argued, for LHC to produce the monopole thermally, the temperature of the fireball must be no less than 119.4 GeV.
+And if the temperature of the fireball higher than this, the above conclusion becomes valid.
+But is this really so?
+How can we guarantee this?
+This is a non-trivial question.
+A straightforward way to answer this question is to measure the temperature of the p-p firebal at LHC.
+And actually the ALICE group has measured the temperature of the p-p fireball of the center of mass energy 2.76 TeV at LHC, and found that the temperature is around 297 MeV, which is roughly $10^{-4}$ times smaller than the center of mass energy of the p-p fireball [27].
+The ALICE result was unexpected, but clearly shows that the temperature of LHC p-p fireball is much less than the temperature to produce the monopole thermally, even with the future FCC energy upgrade.
+This tells that here again, the thermal production of the electroweak monopole at LHC is practically impossible, although it is logically possible.
+This teaches us an important lesson.
+If the monopole production mechanism at LHC is the thermal production, it is vertually impossible for LHC to produce the monopole, simply because the temperature of the p-p fireball is too low.
+In this case the only way to detect the electroweak monopole is to look for the remnant monopoles produced in the earluy universe.
+This also tells that the only probable way for LHC to prooduce the monopole is by the Drell-Yan process and/or the Schwinger mechanism.
+Obviously, the present 14 TeV LHC could produce the monopole if the mass becomes less than 7 TeV, if the monopole production mechanism becomes the Drell-Yan process and/or the Schwinger mechanism.
+But as we have pointed out, in this case there is a possibility for LHC to produce the monopole in the form of the monopolium bound states, even when the monopole mass becomes 11 TeV.
+This is because the monopolium binding energy could reduse the monopolium mass below 14 TeV even when the mass of the monopole-antimonopole pair becomes 22 TeV.
+This necessiate us to study the monopolium bound state, which we discuss in the following.
 V Electroweak Monopolium: Atomic Model
-Let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
+The monopolium itself has been studied before for various reasons [28, 29, 30].
+For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV [28].
+But here we are interested in the electroweak monopolium made of the Cho-Maison monopole pair.
+To discuss the electroweak monopolium we let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
 $\Big{(}-\frac{1}{2\mu}\nabla^{2}+V\Big{)}\Psi=E~{}\Psi,$
@@
 where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 11.0~{}\text{TeV}$.
-So the Rydburg energy of the monopolium becomes 51.615 TeV.
+So the Rydburg energy of the monopolium becomes 51,615 TeV.
 Moreover, the Bohr radius $R$ of the monopolium is given by
 $R=\frac{\alpha}{\mu}\times\alpha^{2}=\frac{2\alpha^{3}}{M}\simeq\frac{5.6\times~{}10^{-9}}{M_{W}}$
-$\simeq 1.4\times 10^{-24}~{}\mbox{cm},$ (51)
+$\simeq 1.4\times 10^{-24}cm,$ (51)
 Notice, however, that these results are totally unreasonable.
-First, the above Rydburg energy is too big (roughly 2,346 times bigger than two monopole masses), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium.
+First, the above Rydburg energy is too big (roughly 2,346 times bigger than the monopole-antimonopole mass), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium.
 Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size.
-This could mean the followings.
+This could mean the following.
 First, the monopolium quickly annihilates itself, long before it reaches to the ground state.
@@
 In this case we have to find a more realistic model of the monopolium.
-The reason for the above results originates from two facts.
-First, the magnetic coupling of the monopole is too strong.
-Second, the monopole mass is too big.
-We could make the atomic model more realistic by improving on these points.
-For example, we could modify the magnetic Coulombic potential at short distances for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$.
-Moreover, we could make the monopole mass less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
-On the other hand, the above exercise has one positive side, because it implies that the binding energy of the electroweak monopolium could be big, of the order of TeV.
-This could make the monopolium mass considerably lighter than $2M$, so that we could have the monopolium of mass less than 14 TeV even when the mass of monopole-antimonopole pair is heavier than 14 TeV.
-With these modifications, we could have a more realistic model of the electroweak monopolium that we discuss now.
-VI Electroweak Monopolium: A Realistic Model
-To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, because the monopole (and antimonopole) has the size of $1/M_{W}$.
-This means that at short distances the singularity of the magnetic potential in (49) should be regularized.
+The reason for the above results originates from the fact that the magnetic coupling of the monopole is too strong, which makes the monipolium size too small.
+We could make the atomic model of the monopolium more realistic by improving on this point.
+For example, we could modify the magnetic Coulombic potential at short distance for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$.
+On the other hand the above exercise has one positive side, because it implies that the binding energy of the electroweak monopolium could be big, of the order of TeV.
+This could make the monopolium mass considerably less than $2M$, so that we could have the monopolium of mass lighter than 14 TeV even when the mass of monopole-antimonopole pair is heavier than 14 TeV, with a more realistic monopolium potential.
+VI A Realistic Electroweak Monopolium
+To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, considering the fact that the monopole (and antiminopole) has the size of $1/M_{W}$.
+This means that at short distance the singularity of the magnetic potential in (49) should be regularized.
 So we modify the potential $V$ at short distance with a harmonic oscillator potential smoothly connected at $r_{0}$, and let
-$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ \\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
+$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ {}\hfil\\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},{}{}{}{}{}{}{}{}&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
 where $r_{0}$ is of the order of $1/M_{W}$. The potential is shown in Fig. 6 in the unit of $r_{0}$.
-Figure 6: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium is shown as a function of $r/r_{0}$.
+Figure 6: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium shown as a function of $r/r_{0}$.
 The significance of this potential is that the coupling constant $k=2\pi/e^{2}r_{0}$ of the hamornic oscillator potential is automatically fixed by the smoothness of the potential at $r_{0}$.
@@
 $R_{\rm I}=x^{l+1}\exp\Big{(}-\frac{x^{2}}{2x_{0}^{3/2}}\Big{)}w(x^{2}/x_{0}^{3/2}),$
-$a_{\rm I}=\frac{2l+3+x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4},~{}~{}~{}~{}b_{\rm I}=l+3/2,$
+$a_{\rm I}=\frac{2l+3+x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4},~{}~{}~{}~{}b_{\rm I}=l+3/2,$ (60)
+and
 $R_{\rm II}=x^{l+1}\exp\big{(}-\sqrt{\epsilon}x\big{)}w(2\sqrt{\epsilon}x),$
@@
 As is well-known, the Kummer’s equation has two independent solutions, the confluent hypergeometric function $M(a,b,z)\equiv\,_{1}F_{1}\bigl{(}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{|}z\bigr{)}$ and the Tricomi’s function $U(a,b,z)$.
+Figure 7: The radial monopolium wave function R(x) for $n=1,2,3,4,5$ shown in (62), (64), and (65).
 In the region I ($x<x_{0}$), we need to satisfy the regularity condition at the origin $x=0$, which force us to exclude $U(a,b,z)$ as it has a singularity at $z=0$ in contrast to $M(a,b,z)$, which is an entire function in $z$.
@@
 $U(a,b,x)\sim x^{-a}\,_{2}F_{0}\bigl{(}\begin{matrix}a&a-b+1\\ &\end{matrix}\bigr{|}-1/x\bigr{)},$ (63)
-as $x$ goes to infinity. Hence, the general solution in the second region has a following form,
+as $x$ goes to infinity. Hence, the general solution in the second region has the following form,
 $R_{\rm II}=N_{\rm II}\,x^{l+1}\exp\big{(}-\sqrt{\epsilon}x\big{)}$
@@
 where $N_{\rm II}$ is again an arbitrary constant.
-Figure 7: The radial monopolium wave function R(x) for $n=1,2,3,4,5$ shown in (63), (69), and (8).
+Figure 8: The energy spectrum $\sqrt{\epsilon}$ of the monopolium with $l=0$ fixed by (65), as function of $\sqrt{x_{0}}$.
+The dashed lines represent harmonic oscillator energy levels given by (67).
+The horizontal dotted lines are the energy levels of the hydrogen atom given by (66).
 The remaining task is to formulate sewing condition that ensures $C_{1}$ continuity of the wavefunction at the point $x=x_{0}$.
@@
 The other condition $R_{\rm I}^{\prime}(x_{0})=R_{\rm II}^{\prime}(x_{0})$ then boils down to a transcendental equation for the binding energy,
-$\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{3+2l}\frac{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{7+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{5}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{3}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}$ (74)
-$+2\sqrt{x_{0}}\big{(}(1+l)\sqrt{\epsilon}-1\big{)}\frac{U\Big{(}\begin{matrix}2+l-\epsilon^{-1/2}\\ 3+2l\end{matrix}\Big{|}2x_{0}\epsilon^{1/2}\Big{)}}{U\Big{(}\begin{matrix}1+l-\epsilon^{-1/2}\\ 2+2l\end{matrix}\Bigr{|}2x_{0}\epsilon^{1/2}\Big{)}}$ (79)
+$\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{3+2l}\frac{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{7+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{5}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{3}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}$
+$+2\sqrt{x_{0}}\big{(}(1+l)\sqrt{\epsilon}-1\big{)}\frac{U\Big{(}\begin{matrix}2+l-\epsilon^{-1/2}\\ 3+2l\end{matrix}\Big{|}2x_{0}\epsilon^{1/2}\Big{)}}{U\Big{(}\begin{matrix}1+l-\epsilon^{-1/2}\\ 2+2l\end{matrix}\Bigr{|}2x_{0}\epsilon^{1/2}\Big{)}}$
 $+\sqrt{\epsilon x_{0}}=1.$ (65)
 With the two conditions we have the smooth radial wave function of the monopolium. This is shown in Fig. 7.
-Figure 8: The contours where Eq.
-(8) holds as function of $\sqrt{x_{0}}$ and $\sqrt{\epsilon}$ with $l=0$.
-The dashed lines represent harmonic oscillator levels given by (82).
-The horizontal dotted lines are the levels of the hydrogen atom, i.e., $\epsilon=1/n^{2}$, with $n=1,2,\ldots$
 Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (8).
 Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig.
-8 for $l=0$.
+8 (for $l=0$).
 We see that as $x_{0}$ goes to zero, the binding energies approaches the values given by the atomic model
@@
 This confirms that the monopolium bound state with the modified potential (6) can give us the desired binding energy which could help LHC to produce the bound state even when the monopole mass becomes bigger than 7 TeV.
-Notice that in the above analysis we have assumed that the monopole mass is 11 TeV.
-However, at LHC the monopole mass at $T_{min}^{LHC}$ becomes 5.3 TeV, so that the monopolium at LHC is expected to be the monopole-antimonopole bound state with mass 10.6 TeV.
-This will change the energy spectrum and the Bohr radius of the monopolium.
-For example, $E_{0}$ in (84) should change to $E_{0}\simeq-8.1~{}\text{TeV}$, which reduces the monopolium mass to 2.5 TeV at LHC.
-It should be mentioned that a different model of the monopolium which assumes an exponential repulsion at short distance given by the potential
+Notice that in the above analysis we have assumed the monopole mass to be 11.0 TeV to have the monopolium mass 5.2 TeV.
+And (69) tells that the monopolium mass becomes 3.3 TeV when the monopole mass becomes 7 TeV.
+Perhaps more importantly, this tells that the maximum mass of the monopole that the present 14 TeV LHC can produce the monopolium of mass 14 TeV becomes 29.8 TeV.
+This strongly implies that the present LHC could produce the electroweak monopole in the form of the monopolium bound states for any reasonable monopole mass (as far as the mass remains less than 29.8 TeV), if the monopole production mechanisn becomes the Drell-Yan process.
+Of course, the monopolium mass at LHC may not turn out to be the same as the above prediction, because we do not know how realistic our potential (55) is.
+In this connection it should be mentioned that a different model of the monopolium which assumes an exponential repulsion at short distance given by the potential
 $V=-\frac{4\pi}{e^{2}}\frac{1}{r}\Big{(}1-\exp(-r/2r_{0})\Big{)},$ (71)
@@
 In particular, in both cases the monopolium masses are slightly below the half of the monopole’s mass and the radiuses are of the order of $10^{-18}cm$.
-This implies that both could be used as a realistic model of the monopolium.
+This implies that our results discussed above are reliable.
 VII Discussion
@@
 Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it.
-But the problem is that the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV.
+But the problem is that intuitively the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV.
 And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, LHC cannot produce it.
-In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition.
-Our work in this paper tells that, if LHC produces the monopole by the topological thermal fluctuation of the Higgs vacuum, LHC could circumbent this energy constraint and produce the electroweak monopole even when the monopole mass becomes 11.0 TeV.
-This is because in this thermal production the monopole mass depends on the temperature at production.
-In this case the mass of the monopoles produced right after the electroweak phase transition (the baby monopole mass) could be as low as 1.0 TeV, even when the monopole mass at zero temperature (the adolescent monopole mass) becomes 11.0 TeV.
-So, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum (the same mechanism as the monopole production in the early universe), the present 14 TeV LHC could produce the electroweak monopole, and thus MoEDAL could detect it.
-Our results suggest that LHC could produce the infant monopoles of mass around 5.3 TeV in average (assuming the adolescent monopole mass is 11 TeV).
-This is really remarkable, because this implies that there is a finite possibility that the MoEDAL detector at LHC could actually detect the infant electroweak monopoles, even though it is not likely to detect the adolescent (i.e., grown-up) monopoles.
-It should be emphasized that this is a unique feature of the finite temperature physics which is absent in the zero temperature physics.
-In particular, this is impossible if the monopole production mechanism at LHC is the Drell-Yan or Schwinger mechanism, where only the adolescent monopole enters.
-From this we can definitely say that the detection of the infant monopole at LHC should be interpreted as the confirmation that the electroweak monopole production mechanism is indeed topological.
-Another important result of our work is that there is a real possibility for LHC to produce the monopoles in the form of the monopolia.
-When LHC produces the monopoles, it should produce them in pairs.
-But since they have strong magnetic attraction some of them could form bound states, although many of the monopole-antimonopole pairs could quickly annihilate.
-An interesting feature of this monopolium is that the mass could be considerably lighter than the mass of the monopole-antimonopole pair, because the monopolium binding energy could be of the order of ten TeV.
-In fact, our study in this paper suggests that the binding energy of the ground state could be as much as 16.7 TeV, so that the monopole-antimonopole pair of the mass 22.0 TeV could form a bound state of the mass 5.3 TeV.
-This provides another way for LHC to circumbent the energy constraint and produce the monopole pairs.
-This tells that there are two ways for LHC to produce the monopole pairs, the monopole and antimonopole as independent particles or the monopole-antimonopole pair together as a monopolium bound state.
-So MoEDAL and ATLAS could try to detect the monopole (and antimonopole) separately, or together as a monopolium.
-Detecting the monopole (and antimonopole) should be straightforward in principle, but detecting the monopolium could be tricky.
-This is because the monopolium has zero magnetic charge, so that we have to look for the magnetically neutral particles of mass of few TeV.
-Moreover, the monopole and antimonopole inside the monopolium could easily annihilate to photons and electron-positron pairs.
-In this case we may have to look for the decay modes of the monopolium.
-One might wonder if we could also have the infant monopoles in cosmology produced in the early universe.
-This may not be possible, because the monopoles in the early universe are in contact with the hot thermal radiation bath which feeds up the energy and allows them to grow up.
-This means that the remnant monopoles from the early universe that we could detect at present are expected to be the adolescent monopoles.
-This implies that the quenched infant monopole at LHC is not a generic phenomenon but an outcome of the limited energy and size of the p-p fireball at LHC.
-Admittedly the idea of the infant monopole at LHC sounds strange and appears hard to swallow.
-We emphasize, however, that this idea actually comes from the fact that in the electroweak phase transition the W (and Higg) boson mass becomes temperature dependent.
-Once we accept this, the idea of the infant monopole follows naturally.
-Moreover, it appears that (as far as we know) logically it is difficult to exclude this possibility.
-Ultimately, of course, only experiment could tell whether this idea is realistic or not.
-We hope that our results in this paper could help MoEDAL to detect the electroweak monopole at LHC.
+In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition [15, 18].
+In this paper we have shown that the question if the present LHC could produce the monopole or not critically depends on the monopole production mechanism at LHC.
+Our work in this paper tells the followings.
+First, if the monopole production mechanism at LHC is the topological thermal fluctuation of the Higgs vacuum, there is practically no hope that it could produce the monopole, simply because the temperature of the p-p fireball is too low to produce the monopole thermally.
+This is unexpected, but if the recent ALICE measurement of the temperature of the p-p fireball is trustable, this conclusion is unavoidable.
+Second, if the monopole production mechanism at LHC is the Drell-Yan process, the present 14 TeV LHC could produce the monopole if the monopole mass becomes less than 7 TeV.
+This was expected.
+What is new is that, more probably LHC could produce the monopole in the form of the monopolium bound states for any reasonable mass of the monopole, as far as the mass does not exceed 29,7 TeV.
+This is because the monopolium binding energy greatly reduces the mass of the monopolium.
+This tells that the present LHC could circumbent the energy constraint and produce the electroweak monopole even when the monopole mass becomes bigger than 7 TeV. This is remarkable.
+Of course, with the Schwinger mechanism at LHC we can have the same conclusion, if the p-p fireball could create strong enough magnetic background which can actually produce the monopole-antimonopole pairs.
+But it is by no means clear how the p-p fireball at LHC could generate such a strong magnetic background.
+To summarize, our results tells that a most probable way for LHC to produce the monopole is the Drell-Yan process, and the 14 TeV LHC could more likely produce the monopole in the form of the monopolium bound states than as individual monopoles.
+In this case the most realistic way to detect the monopole at LHC is to detect the two photon decay modes of the monopoliums.
+And in principle ATLAS and CMS could do that.
+But there is the possibility that the LHC (with the next energy upgrading) may not produce the monopole signal at all.
+This could mean that the monopole production mechanism at LHC is probably not the Drell-Yan process.
+In this case the only way to detect the electroweak monopole could be to look for the remnant monopoles produced during the early universe, with the “cosmic” MoEDAL.
+This makes the detection of the remnant monopoles produced during the early universe very important.
 ACKNOWLEDGEMENT
@@
 YMC is supported in part by the National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University, Korea.
-The authors thank their supports.

### v5

@@
 Abstract
-The monopole production at LHC crucially depends on the monopole production mechanism.
-We show that, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum as the early universe did, it is practically impossible for LHC to produce the monopole, even with the future FCC energy upgrade.
-This is because the temperature of the p-p fireball is simply too low to produce the monopole.
-But if the monopole production mechanism is the Drell-Yan and/or Schwinger mechanism, the 14 TeV LHC could produce the monopole in the form of the monopolium of mass around 5,7 TeV even when the monopole mass becomes 11 TeV.
-Or, it could produce the monopole when the mass is less than 7 TeV.
-Our result tells that, if the monopole production mechanism at LHC becomes the thermal fluctuation, searching for the remnant monopoles produced in the early universe could be the only way to detect the electroweak monopole.
+It is argued that the monopole production at LHC crucially depends on the monopole production mechanism.
+We show that, if the monopole production mechanism at LHC becomesthe thermal fluctuation of the Higgs vacuum as the early universe did, it is practically impossible for LHC to produce the monopole.
+This is because the temperature of the p-p fireball is simply too low to generate the thermal fluctuation necessary for the monopole production.
+But if the monopole production mechanism becomes the Drell-Yan and/or Schwinger mechanism, the 14 TeV LHC could produce the monopole only when the mass is less than 7 TeV.
+To circumbent this energy constraint, we propose a new monopole production mechanism at LHC, the monopole production in the form of the monopolium bound state.
+We show that LHC could produce the monopole in the form of the monopolium of mass less than 14 TeV, even when the monopole mass becomes considerably larger than 11 TeV.
+In this case we can prove the existence of the electroweak monopole at present LHC by detecting the decay modes of the monopolium.
+This implies that the monopole production in the form of the monopolium could be the most probable way to detect the electroweak monopole at LHC.
 We discuss the physical implications of our result.
-electroweak monopole, monopole production mechanism at LHC, Drell-Yan process, Schwiner mechanism, thermal fluctuation of Higgs vacuum, Ginzburg temperature, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
+electroweak monopole, monopole production mechanism at LHC, Drell-Yan process, Schwiner mechanism, monopolre production by thermal fluctuation of Higgs vacuum, monopokle production in the form of the monopolium, Ginzburg temperature, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
 I Introduction
@@
 After the Dirac monopole we have had the Wu-Yang monopole, the ’tHooft-Polyakov monopole, and the grand unification monopole [3, 4, 5].
-But the electroweak (Cho-Maison) monopole, stands out as the most realistic monopole that could exist in nature and could actually be detected [6, 7].
+But the electroweak (“Cho-Maison”) monopole stands out as the most realistic monopole that could exist in nature and could actually be detected [6, 7].
 Indeed the Dirac monopole in electrodynamics should transform to the electroweak monopole after the unification of the electromagnetic and weak interactions, and the Wu-Yang monopole in QCD is supposed to make the monopole condensation to confine the color.
 Moreover, the ’tHooft-Polyakov monopole exists only in an hypothetical theory, and the grand unification monopole which could have been amply produced at the grand unification scale in the early universe probably has become completely irrelevant at present universe after the inflation.
-This makes the experimental confirmation of the electroweak monopole one of the most urgent issues in the standard model after the discovery of the Higgs particle.
+This makes the experimental confirmation of the electroweak monopole one of the most urgent issues in the standard model after the discovery of the Higgs particle [8, 9, 10].
 In fact the detection of this monopole, not the Higgs particle, should be regarded as the final (and topological) test of the standard model.
-For this reason the newest MoEDAL detector at LHC is actively searching for the monopole [8, 9, 10, 11].
-To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [6, 7, 12, 13, 14, 15, 16].
+For this reason the MoEDAL and ATLAS detectors at LHC are actively searching for the monopole [11, 12, 13, 14].
+To detect the electroweak monopole at LHC, we need to remember the basic facts about the monopole [6, 7, 8, 15, 16, 17, 18].
 First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’tHooft-Polyakov monopoles.
@@
 This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the magnetic coupling makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV [19].
-Despite this, the size of the monopole is set by the W boson mass, because the monopole solution has the Higgs and W boson dressing which fixes the size by the W boson mass.
+Despite this, the size of the monopole is set by the W boson mass, because the monopole solution has the W (and Higgs) boson dressing which fixes the size by the W boson mass.
 Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it.
 However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV.
-This is problematic, because the monopole mass could turn out to be bigger than this.
+This is problematic, because the monopole mass could turn out to be bigger than this, around 11 TeV.
 If this is so, we may have little chance to detect the monopole at LHC, and may have to try to detect the remnant monopoles at present universe produced in the early universe [8, 20].
-However, two observations could make the monopole production at LHC possible even when the mass becomes over 7 TeV.
+However, two observations could make the monopole production at LHC possible even when the mass becomes heavier than 7 TeV.
 First, the monopole has to be created in pairs at LHC.
-This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole.
+This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and anti-monopole pair.
 In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
 LHC could produce the monopole with this mechanism because the fireball of the p-p collision could reproduce the early universe.
-A unique feature of this topological thermal fluctuation mechanism is that in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [15, 18].
-In this case there is a possibility that LHC could produce the electroweak monopole even when the zero temperature mass becomes heavier than 7 TeV.
+A unique feature of this topological thermal fluctuation mechanism is that in this production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [8, 20].
+In this case there is a possibility that LHC could produce the baby electroweak monopole even when the zero temperature mass becomes heavier than 7 TeV.
 The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV.
-We show that, if the monopole production mechanism at LHC is the Drell-Yan process (and/or Schwinger mechanism), the present 14 TeV LHC could produce the monopole if the mass is less than 7 TeV.
-Or, it could produce the monopole in the form of the monopolium bound states of mass around 5.7 TeV even when the monopole mass becomes 11 TeV.
+We confirm that, if the monopole production mechanism at LHC is the Drell-Yan process (and/or Schwinger mechanism), the present 14 TeV LHC could produce the monopole only if the mass is less than 7 TeV.
+However, if the monopole production mechanism at LHC becomes the thermal fluctuation of the Higgs vacuum, it has practically no chance to produce the monopole (even with the future FCC).
+This is because the temperature of the p-p fireball at LHC simply becomes too low to produce the monopole thermally.
+We propose a new monopole production mechanism at LHC, the monopole production in the form of the monopolium, which could allow LHC to produce the electroweak monopole even when the monopole mass becomes larger than 7 TeV.
+In specific, we show that LHC could produce the monopolium bound states of mass around 5.7 TeV for any reasonable monopole mass, as far as the mass does not exceed 29.8 TeV.
 This is because the binding energy of the monopoliun could reduce the monopolium mass by 16.3 TeV.
-Indeed we argue that the LHC can produce the monopolium for any reasonable monopole mass, as far as the mass does not exceed 29.8 TeV.
-However, if the monopole production mechanism at LHC becomes the thermal fluctuation of the Higgs vacuum, it has practically no chance to produce the monopole even with the future FCC energy upgrade.
-This is because the temperature of the p-p fireball at LHC simply becomes too low to produce the monopole thermally.
-This makes the monopolium bound state a most probable signal for the monopole production at LHC.
+This makes the monopolium bound state a most probable signal for the electroweak monopole production at LHC.
 The paper is organized as follows.
-In Section II we discussthe electroweak monopole production mechanism at LHC.
+In Section II we discuss two electroweak monopole production mechanisms at LHC, Drell-Yan process and Schwinger mechanism.
 In Section III we review the monopole production in the electroweak phase transition in the early universe.
 In Section IV we discuss if LHC could produce the monopole in a similar manner, and argue that this is practically impossible because the temperature of the p-p fireball is too low.
-In Section V we discuss the naive Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium.
-In Section VI we discuss a more realistic electroweak monopolium that could be produced at LHC.
+In Section V we discuss the naive Bohr model of thelectroweak monopolium and show that this model is unrealistic to describe an electroweak monopolium.
+In Section VI we discuss a more realistic electroweak monopole production mechanism, the monopole production via monopolium, and argue that such monopolium bound state could be produced at LHC.
 In Section VII we discuss the physical impications of our results.
-II Electroweak Monopole Production Mechanism at LHC
+II Electroweak Monopole Production Mechanism at LHC: Drell-Yan Process and Schwinger Mechanism
 Consider the (bosonic sector of the) Weinberg-Salam Lagrangian,
@@
 $|{\cal D}_{\mu}\xi|^{2}=\frac{1}{4}Z_{\mu}^{2}+\frac{g^{2}}{2}W_{\mu}^{*}W_{\mu}.$ (18)
-Notice that here we have defined $A_{\mu}^{(em)}$ and $Z_{\mu}$ gauge independently, without choosing the unitary gauge.
+Notice that here we have defined $A_{\mu}^{(em)}$ and $Z_{\mu}$ gauge independently, without any reference to the unitary gauge.
 Figure 1: The Feynman diagrams of the popular monopole production mechanism at LHC given by Drell-Yan and two-photon fusion process. Here the proton pair could also be interpreted as the quark pair.
@@
 Figure 2: The topological monopole production mechanism in the early universe or at LHC induced by the thermal fluctuation of the Higgs vacuum.
-The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field.
+The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $CP^{1}$ field.
 It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [6, 7].
 Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV.
-Intuitively we could argue the mass to be $1/\alpha$ times bigger than the W boson mass, around 11 TeV [17].
-This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$).
+Intuitively we could argue the mass to be $1/\alpha$ times bigger than the W boson mass, around 11 TeV.
+This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$) [19].
 We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [15, 16, 17, 8, 18].
@@
 But in the Schwinger mechanism a strong magnetic field is supposed to create the monopole-antimonopole pair, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [13, 21].
-The justification for this comes from the asumption that a theory of monopole must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$.
+The justification for this comes from the asumption that a theory of monopole must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to (g,e)(g,-e).
 The Drell-Yan process is shown graphically in Fig.
 1.
-But in the topological monopole production, the thermal fluctuation of the Higgs vacuum induces the change of topology and produce the monopole, which is exactly the electroweak monopole production mechanism in the early universe [15].
-In this picture the monopoles are produced after the phase transition and stops at the Ginzburg temperature [15].
+But in the topological monopole production, the thermal fluctuation of the Higgs vacuum induces the change of topology and produce the monopole, which is exactly the monopole production mechanism in the early universe [8].
+In this picture the monopoles are produced after the phase transition and stops at the Ginzburg temperature.
 The topological monopole production mecanism is shown graphically in Fig.
 2 for comparison.
-Notice that, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process.
-But the Weinberg-Salam Lagrangian perturbatively has no interaction which can be described by the above Feynman diagram.
+Notice that, in the Drell-Yan process the monopole pair is produced via the photon solely by the electromagnetic process.
+But the Weinberg-Salam Lagrangian has no interaction which can be described by the above Feynman diagram perturbatively.
 Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background.
-Most importantly, we need the change of topology to produce the monopole, but this topological nature of the monopole has no role at all in these mechanisms.
+So these monopole production mechanisms are possible beyond (not within) the standard model.
+More importantly, these mechanisms do not take into account the change of topology necessary to produce the monopole, so that the topological nature of the monopole is completely neglected in these mechanisms.
 In comparison the topology plays an essential role in the thermal production of the monopole, because here the thermal fluctuation of the Higgs vacuum is precisely what we need to induce the change of topology.
 As a result the monopole production in this mechanism is not controlled by any fundamental interaction or fundamental constant.
-Perhaps more importantly, in this thermal fluctuation the monopole mass depends on the temperature at creation, the monopole mass at creation could be considerably smaller than the monopole mass at zero temperature.
+Perhaps more importantly, in this thermal fluctuation the monopole mass depends on the temperature at creation because the vacuum value of the Higgs field which determines the monopole mass does.
+This means that the monopole mass at creation could be considerably smaller than the monopole mass at zero temperature.
 In this case we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass.
 So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” (or “infant”) monopole mass and the zero temperature mass as the “adolescent” monopole mass.
-It should be emphasized that the above monopole production mechanisms at LHC are the logical possibilities, not the realistic possibilities.
+It should be emphasized that the above monopole production mechanisms at LHC are theoretical (i.e., logical) but not realistic possibilities.
 For example, the Schwinger mechanism is certainly a logical possibility.
 However, it is not clear at all if (and how) the p-p fireball at LHC could create a strong magnetic background which can actually produce the monopole-antimonopole pair.
-If so, what is the the monopole production mechanism at LHC?
+If so, what is the monopole production mechanism at LHC?
 We do not know yet.
 On the other hand, it has often been claimed that LHC could reproduce the early universe.
-If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production in the early universe.
-In the following we discuss if LHC could produce the electroweak monopole thermally.
-III Monopole Production in Early Universe: A Review
+If this is true, the monopole production mechanism at LHC could be thermal, just like the electroweak monopole production in the early universe.
+To discuss if this is true, we have to understand the monopole production mechanism in the early universe.
+III Electroweak Monopole Production in the Early Universe: A Review
 It has generally been believed that the monopole production in the early universe critically depends on the type of the phase transition.
 In the first order phase transition the vacuum bubble collisions in the unstable vacuum are supposed to create the monopoles through the quantum tunneling to the stable vacuum during the phase transition, so that the monopole production is supposed to be suppressed exponentially by the vacuum tunneling [23].
-On the other hand the monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism with no exponential suppression [22, 23].
+On the other hand the monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism which has no such exponential suppression [24, 25].
 We emphasize, however, that this popular view may have a critical defect [8].
@@
 This tells that what is important in the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
-To amplify this point we start from the temperature dependent effective action of the standard model which describes the electroweak monopole production in the early universe [15, 25, 26, 18]
+To amplify this point we start from the temperature dependent effective action of the standard model which describes the electroweak phase transition and the monopole production in the early universe [8, 27, 28, 20]
 $V_{eff}(\rho)=V_{0}(\rho)-\frac{C_{1}}{12\pi}\rho^{3}~T+\frac{C_{2}}{2}\rho^{2}~T^{2}-\frac{\pi^{2}}{90}N_{*}T^{4}$
@@
 So for all practical purposes we could treat the electroweak phase transition as a second order phase transition.
-Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the black line represents $T_{G}$.
+Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition.
+Here the red and blue curve represents $\xi^{3}\Delta F$ and the temperature of the universe, and the black line represents $T_{G}$.
 The monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism, so that the monopole production start from $T_{c}$.
@@
 $\rho_{+}(T_{i})\simeq 188.4~\text{GeV}.$ (30)
-To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [kolb]
+To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [29]
 $t=\Big{(}\frac{90}{32\pi GN_{*}(T)}\Big{)}^{1/2}~\frac{1}{T^{2}}.$ (31)
@@
 $\bar{M}_{H}^{2}=\frac{d^{2}V_{eff}}{d\rho^{2}}\Big{|}_{\rho_{min}}$
-=$=\left\{\begin{array}[]{ll}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T>T_{c},\\ \Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
+$=\left\{\begin{array}[]{ll}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~\dfrac{M_{H}^{2}}{2},&~~~T>T_{c},\\ \Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~\dfrac{M_{H}^{2}}{2},&~~~T\leq T_{c},\end{array}\right$.
+(35)
 and the W-boson mass which determines the monopole mass $M\simeq\bar{M}_{W}/\alpha$,
@@
 The temperature dependent Higgs and W-boson masses are shown in Fig. 5.
-$T$ $\rho_{+}(T)$ $\bar{M}_{H}(T)$ $\bar{M}_{W}(T)$ $\bar{M}(T)$
-$T_{1}$    $146.74$ $16.3$ $5.9$ 00 00
-$T_{c}$ $146.70$ $21.7$ $5.5$ $7.1$ $971.7$
-$T_{2}$ $146.42$ $32.5$ $11.7$ $0.61$ 454.81454.81\,454.81 454.8
-$T_{i}^{LHC}$ $132.89$ $119.2$ $56.8$ $38.9$ 5 331.55331.55\,331.55 331.5
-$T_{min}^{LHC}$ $119.36$ $156.5$ $76.2$ $51.1$ 7 000.07000.07\,000.07 000.0
-$T_{i}$ $101.96$ $188.4$ $92.9$ $61.5$ 8 428.28428.28\,428.28 428.2
-$T_{G}$ $57.49$ $232.9$ $116.9$ $76.0$ 10 419.610419.610\,419.610 419.6
-00 $0.00$       $246.2$       $125.3$       $80.4$ 11 014.511014.511\,014.511 014.5
-Table 1: The values of $\rho_{+}$, $\bar{M}_{H}$, $\bar{M}_{W}$, and the expected monopole mass $\bar{M}=\bar{M}_{W}/\alpha$ at various temperatures. All numbers are in GeV.
 Notice that the Higgs boson acquires the minimum mass $\bar{M}_{H}\simeq 5.5~{\rm GeV}$ at $T_{c}$ and approaches to the zero temperature mass 125.3 GeV as the universe cools down. Moreover, we have
 $\bar{M}_{H}(T_{G})\simeq 116.9~{\rm GeV}.$ (39)
-This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature.
+This confirms that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature.
 But the W-boson which is massless above $T_{c}$ (before the symmetry breaking) becomes massive at $T_{c}$, and we have
@@
 This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~\text{TeV}$).
-This confirms that the baby monopole mass near $T_{c}$ could be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value.
+This tells that the baby monopole mass near $T_{c}$ could be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value.
 This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV.
 In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
+$T$ $\rho_{+}(T)$ $\bar{M}_{H}(T)$ $\bar{M}_{W}(T)$ $\bar{M}(T)$
+$T_{1}$ $146.74$ $16.3$ $5.9$ 0 0
+$T_{c}$ $146.70$ $21.7$ $5.5$ $7.1$ $971.7$
+$T_{2}$ $146.42$ $32.5$ $11.7$ $10.6$ 1 454.81\,454.8
+$T_{i}^{LHC}$ $132.89$ $119.2$ $56.8$ $38.9$ 5 331.55\,331.5
+$T_{min}^{LHC}$ $119.36$ $156.5$ $76.2$ $51.1$ 7 000.07\,000.0
+$T_{i}$ $101.96$ $188.4$ $92.9$ $61.5$ 8 428.28\,428.2
+$T_{G}$ $57.49$ $232.9$ $116.9$ $76.0$ 10 419.610\,419.6
+0 $0.00$     $246.2$     $125.3$     $80.4$ 11 014.511\,014.5
+Table 1: The values of $\rho_{+}$, $\bar{M}_{H}$, $\bar{M}_{W}$, and the expected monopole mass $\bar{M}=\bar{M}_{W}/\alpha$ at various temperatures. All numbers are in GeV.
 This also confirms that what is important for the monopole production in the early universe is the Ginzburg temperature, not the type of the phase transition.
@@
 As far as the Ginzburg temperature becomes lower than $T_{2}$, there is no much difference between the monopole production in the first and second order phase transitions.
-IV Electroweak Monopole Production from p-p Fireball at LHC
-If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it has been asserted, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does.
+IV Topological Monopole Production by thermal fluctuation of Higgs vacuum at LHC
+If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it has often been asserted, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does.
 In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
-However, the 14 TeV LHC can not produce the monopole of mass 10.4 TeV.
+However, the 14 TeV LHC can not produce the monopole of mass 10.4 TeV as it has to produce the monopole in pairs.
 This means that the present LHC can not produce the monopole exactly as the early universe does.
@@
 The effective potential at $T_{min}^{LHC}$ is also shown in Fig.
-3 for comparison.
-If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature.
+3 in green line for comparison.
+If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature around 57.5 GeV.
 In average the monopole production temperature at LHC is given by
@@
 This assumotion, however, may have a critical defect.
-This is because the correlation length is fixed by the electroweak scale but the monopole mass is given by the $1/\alpha$ times bigger than the electroweak scale, so that the energy in one correlation volume may not be enough to make up the monopole mass.
+This is because the correlation length is fixed by the electroweak scale but the monopole mass is given by $1/\alpha$ times bigger than the electroweak scale, so that the energy in one correlation volume may not be enough to make up the monopole mass.
 A natural way to cure this defect is to enlarge the correlation length $\xi=1/\bar{M}_{H}$ to $\bar{\xi}$,
@@
 Unfortunately, the above argument has a critical defect, because here we have implicitly assumed that the temperature of the p-p fireball would be high enough to produce the monopole thermally (assuming that the 14 TeV p-p fireball energy would produce the high temperature necessary for LHC to produce the monopole thermally).
-But this is not obvious, because the temperature and energy of the fireball are two different things.
+However, this is not obvious, because the temperature and energy of the fireball are two different things.
 As we have argued, for LHC to produce the monopole thermally, the temperature of the fireball must be no less than 119.4 GeV.
@@
 But is this really so?
-How can we guarantee this?
 This is a non-trivial question.
 A straightforward way to answer this question is to measure the temperature of the p-p firebal at LHC.
-And actually the ALICE group has measured the temperature of the p-p fireball of the center of mass energy 2.76 TeV at LHC, and found that the temperature is around 297 MeV, which is roughly $10^{-4}$ times smaller than the center of mass energy of the p-p fireball [27].
-The ALICE result was unexpected, but clearly shows that the temperature of LHC p-p fireball is much less than the temperature to produce the monopole thermally, even with the future FCC energy upgrade.
+Fortunately the ALICE group has actually measured the temperature of the p-p fireball of the center of mass energy 2.76 TeV at LHC, and found that the temperature is around 297 MeV, which is roughly $10^{-4}$ times smaller than the center of mass energy of the p-p fireball [30].
+The ALICE result was unexpected, but clearly shows that the temperature of LHC p-p fireball is much less than the temperature needed to produce the monopole thermally, even with the future FCC energy.
 This tells that here again, the thermal production of the electroweak monopole at LHC is practically impossible, although it is logically possible.
 This teaches us an important lesson.
-If the monopole production mechanism at LHC is the thermal production, it is vertually impossible for LHC to produce the monopole, simply because the temperature of the p-p fireball is too low.
+If the monopole production mechanism at LHC is the thermal production, it is virtually impossible for LHC to produce the monopole, simply because the temperature of the p-p fireball is too low.
 In this case the only way to detect the electroweak monopole is to look for the remnant monopoles produced in the earluy universe.
-This also tells that the only probable way for LHC to prooduce the monopole is by the Drell-Yan process and/or the Schwinger mechanism.
-Obviously, the present 14 TeV LHC could produce the monopole if the mass becomes less than 7 TeV, if the monopole production mechanism becomes the Drell-Yan process and/or the Schwinger mechanism.
-But as we have pointed out, in this case there is a possibility for LHC to produce the monopole in the form of the monopolium bound states, even when the monopole mass becomes 11 TeV.
-This is because the monopolium binding energy could reduse the monopolium mass below 14 TeV even when the mass of the monopole-antimonopole pair becomes 22 TeV.
-This necessiate us to study the monopolium bound state, which we discuss in the following.
+This also seems to suggest that the Drell-Yan process and/or the Schwinger mechanism might be the only probable way for LHC to produce the monopole.
+Obviously, the present 14 TeV LHC could produce the monopole when the mass becomes less than 7 TeV, if the monopole production mechanism becomes the Drell-Yan process and/or the Schwinger mechanism.
+On the other hand, there may be another logical possibility for LHC to produce the monopole, the monopole production in the form of the monopolium bound states, as we have pointed out.
+An important feature of this mechanism is that the binding energy of the monopolium could reduce the mass below 14 TeV even when the mass of the monopole-antimonopole pair becomes 22 TeV.
+In this case LHC could produce the monopolium even when the monopole mass becomes 11 TeV or heavier.
+This necessiate us to study the monopolium bound state in more detail, which we discuss in the following.
 V Electroweak Monopolium: Atomic Model
 The monopolium itself has been studied before for various reasons [31, 32, 33].
-For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV [28].
+For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum of the monopole pair to the order of hundred MeV [31].
 But here we are interested in the electroweak monopolium made of the Cho-Maison monopole pair.
@@
 First, the above Rydburg energy is too big (roughly 2,346 times bigger than the monopole-antimonopole mass), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium.
-Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size.
+Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size, which is absurd.
+Figure 6: The monopole production mechanismin the form of the monopolium at LHC.
 This could mean the following.
-First, the monopolium quickly annihilates itself, long before it reaches to the ground state.
+First, the monopolium quickly annihilates itself, long before it decays to the ground state.
 In fact (50) suggests that the monopolium could exist only at highly excited levels, probably with $n\geq 49$, because below this level the monopole-antimonopole mass can not provide the necessary binding energy.
@@
 In this case we have to find a more realistic model of the monopolium.
-The reason for the above results originates from the fact that the magnetic coupling of the monopole is too strong, which makes the monipolium size too small.
+The reason for the above results originates from the fact that the magnetic coupling of the monopole is too strong, which makes the monopolium size too small.
 We could make the atomic model of the monopolium more realistic by improving on this point.
@@
 This could make the monopolium mass considerably less than $2M$, so that we could have the monopolium of mass lighter than 14 TeV even when the mass of monopole-antimonopole pair is heavier than 14 TeV, with a more realistic monopolium potential.
-VI A Realistic Electroweak Monopolium
-To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, considering the fact that the monopole (and antiminopole) has the size of $1/M_{W}$.
+VI Electroweak Monopole Production at LHC via Monopolium
+Now, we are ready to discuss the new monopole production mechanism in the form of the monopolium, which could allow us to prove the existence of the electroweak monopole at LHC even when the monopole mass exceeds 11 TeV.
+This is schimatically shown in Fig.
+6.
+Notice that LHC could produce the monopole with this mechanism, when the monopole production mechanism at LHC becomes the Drell-Yan process and/or the Schwinger mechanism but the energy condition prevent the production of the monopole when the monopole mass becomes bigger than 7 TeV.
+In fact the first stage of this mechanism is exactly like the Drell-Yan process which produces the monopole-antimonopole pair.
+But when the monopole mass exceeds 7 TeV, the pair can not be materialized and could exist only vertually.
+So, the Drell-Yan process becomes “virtual”, and the final state of this mechanism becomes not the monopole-antimonopole pair but the decay modes of the monopolium.
+We could have a similar situation with the Schwinger mechanism.
+Here again the energy condition of LHC could force the monopole-antimonopole pair virtual, and the final state could become the decay modes of the monopolium.
+In this sence this monopole production mechanism could be thought as the virtual Drell-Yan process or the virtual Schwinger mechanism.
+On the other hand, we emphasize that the final state of this monopole production mechnism is totally different from the Drell-Yan process and/or the Schwinger mechanism.
+To show that this mechanism could indeed prove the existence of the electroweak monopole at LHC even when the monopole mass exceeds 11 TeV, we must have a realistic model of the monopolium whose binding enerfy could reduce the monopolum mass below 14 TeV.
+For this we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, considering the fact that the monopole has the size of $1/M_{W}$.
 This means that at short distance the singularity of the magnetic potential in (49) should be regularized.
+Figure 7: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium shown as a function of $r/r_{0}$.
 So we modify the potential $V$ at short distance with a harmonic oscillator potential smoothly connected at $r_{0}$, and let
-$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ {}\hfil\\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},{}{}{}{}{}{}{}{}&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
+$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~~~r<r_{0},\\ \\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},&~~~r\geq r_{0},\end{array}\right$. (55)
 where $r_{0}$ is of the order of $1/M_{W}$. The potential is shown in Fig. 7 in the unit of $r_{0}$.
-Figure 6: The modified magnetic potential (55) (multiplied by $r_{0}$) of the electroweak monopolium shown as a function of $r/r_{0}$.
 The significance of this potential is that the coupling constant $k=2\pi/e^{2}r_{0}$ of the hamornic oscillator potential is automatically fixed by the smoothness of the potential at $r_{0}$.
@@
 With this observation we could solve the Schrödinger equation keeping the reduced mass $\mu$ of the monopole-antimonopole pair arbitrary, allowing $\mu$ to have a small mass.
-The first step is the usual separation of variables for wave functions in the regions $r<r_{0}$ (region I) and $r>r_{0}$ (region II) as
-$\Psi_{\rm{I,II}}(r,\varphi,\theta)\equiv\frac{R_{\rm{I,II}}(r)}{r}Y_{l}^{m}(\varphi,\theta)\,,$ (56)
-where $Y_{l}^{m}(\varphi,\theta)$ are the spherical harmonics. Next, we adopt dimensionless coordinates
+The first step to do this is the usual separation of variables of the wave function in the regions $r<r_{0}$ (the region I) and $r>r_{0}$ (the region II) as
+$\Psi_{\rm{I,II}}(r,\theta,\varphi)\equiv\frac{R_{\rm{I,II}}(r)}{r}Y_{l}^{m}(\theta,\varphi),$ (56)
+where $Y_{l}^{m}(\theta,\varphi)$ are the spherical harmonics. Next, we adopt dimensionless coordinates
 $r=\frac{2\alpha}{M}x\,~~~~r_{0}=\frac{2\alpha}{M}x_{0},~~~~E\equiv-\frac{M}{4\alpha^{2}}\epsilon,$ (57)
@@
 $R_{\rm II}=x^{l+1}\exp\big{(}-\sqrt{\epsilon}x\big{)}w(2\sqrt{\epsilon}x),$
-$a_{\rm II}=l+1-\epsilon^{-1/2},\quad b_{\rm II}=2l+2.$ (61)
+$a_{\rm II}=l+1-\epsilon^{-1/2},~~~b_{\rm II}=2l+2.$ (61)
 As is well-known, the Kummer’s equation has two independent solutions, the confluent hypergeometric function $M(a,b,z)\equiv\,_{1}F_{1}\bigl{(}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{|}z\bigr{)}$ and the Tricomi’s function $U(a,b,z)$.
-Figure 7: The radial monopolium wave function R(x) for $n=1,2,3,4,5$ shown in (62), (64), and (65).
-In the region I ($x<x_{0}$), we need to satisfy the regularity condition at the origin $x=0$, which force us to exclude $U(a,b,z)$ as it has a singularity at $z=0$ in contrast to $M(a,b,z)$, which is an entire function in $z$.
+In the region I ($x<x_{0}$), we need to satisfy the regularity condition at the origin $x=0$, which forces us to exclude $U(a,b,z)$ as it has a singularity at $z=0$ in contrast to $M(a,b,z)$, which is an entire function in $z$.
 Hence, the general solution in the first region that is regular at the origin is given by
 $R_{\rm I}=N_{\rm I}\,x^{l+1}\exp\Big{(}-\frac{x^{2}}{2x_{0}^{3/2}}\Big{)}$
-×$\times_{1}F_{1}\Big{(}\begin{matrix}\frac{l}{2}+\frac{3}{4}+\frac{x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4}\\ l+\frac{3}{2}\end{matrix}\Bigr{|}\tfrac{x^{2}}{x_{0}^{3/2}}\Big{)},$ (62)
-where $N_{\rm I}$ is an arbitrary constant at the moment.
+$\times_{1}F_{1}\Big{(}\begin{matrix}\frac{l}{2}+\frac{3}{4}+\frac{x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4}\\ l+\frac{3}{2}\end{matrix}\Bigr{|}\tfrac{x^{2}}{x_{0}^{3/2}}\Big{)},$ (62)
+where $N_{\rm I}$ is an arbitrary constant at the moment which should be determined later.
 In the region II ($x>x_{0}$), we need to impose boundary condition at $x\to\infty$ so that the resulting wave equation is square-integrable.
-This force us to discard $M(a,b,z)$ that behaves as $M(a,b,z)\sim\Gamma(b)\exp(z)~{}z^{a-b}/\Gamma(a)$ for large $z$ (in a certain wedge of the complex plane).
+This forces us to discard $M(a,b,z)$ that behaves as $M(a,b,z)\sim\Gamma(b)\exp(z)~z^{a-b}/\Gamma(a)$ for large $z$ (in a certain wedge of the complex plane).
 Notice that this asymptotic behaviour holds only if $a$ is not a negative integer (and hence $\Gamma(a)$ blows up), which is exactly the case for the hydrogen atom.
-For our modified potential, however, the spectrum differs and hence $a\not=-k$, where $k$ is a positive integer.
+For our modified potential, however, the spectrum differs and hence $a\neq-k$, where $k$ is a positive integer.
 On the other hand, we have
 $U(a,b,x)\sim x^{-a}\,_{2}F_{0}\bigl{(}\begin{matrix}a&a-b+1\\ &\end{matrix}\bigr{|}-1/x\bigr{)},$ (63)
-as $x$ goes to infinity. Hence, the general solution in the second region has the following form,
+as $x$ goes to infinity. Hence, the general solution in the region II has the following form,
 $R_{\rm II}=N_{\rm II}\,x^{l+1}\exp\big{(}-\sqrt{\epsilon}x\big{)}$
-×$\times U\Big{(}\begin{matrix}l+1-\frac{1}{\sqrt{\epsilon}}\\ 2l+2\end{matrix}\Bigr{|}2\sqrt{\epsilon}x\Big{)},$ (64)
-where $N_{\rm II}$ is again an arbitrary constant.
-Figure 8: The energy spectrum $\sqrt{\epsilon}$ of the monopolium with $l=0$ fixed by (65), as function of $\sqrt{x_{0}}$.
-The dashed lines represent harmonic oscillator energy levels given by (67).
-The horizontal dotted lines are the energy levels of the hydrogen atom given by (66).
+$\times U\Big{(}\begin{matrix}l+1-\frac{1}{\sqrt{\epsilon}}\\ 2l+2\end{matrix}\Bigr{|}2\sqrt{\epsilon}x\Big{)},$ (64)
+where $N_{\rm II}$ is again an arbitrary constant for the moment.
 The remaining task is to formulate sewing condition that ensures $C_{1}$ continuity of the wavefunction at the point $x=x_{0}$.
-The condition that $R_{\rm I}(x_{0})=R_{\rm II}(x_{0})$ fixes, say, $N_{\rm II}$ in terms of $N_{\rm I}$, which can be in turn fixed by square integrability condition.
-The other condition $R_{\rm I}^{\prime}(x_{0})=R_{\rm II}^{\prime}(x_{0})$ then boils down to a transcendental equation for the binding energy,
+The condition that $R_{\rm I}(x_{0})=R_{\rm II}(x_{0})$ fixes $N_{\rm II}$ in terms of $N_{\rm I}$, which can be in turn fixed by square integrability condition.
+The other condition $R_{\rm I}^{\prime}(x_{0})=R_{\rm II}^{\prime}(x_{0})$ then boils down to the following transcendental equation for the binding energy,
 $\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{3+2l}\frac{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{7+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{5}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}{{}_{1}F_{1}\Big{(}\begin{matrix}\frac{3+2l-3x_{0}^{1/2}+x_{0}^{3/2}\epsilon}{4}\\ \frac{3}{2}+l\end{matrix}\Bigr{|}x_{0}^{1/2}\Big{)}}$
@@
 $+\sqrt{\epsilon x_{0}}=1.$ (65)
-With the two conditions we have the smooth radial wave function of the monopolium. This is shown in Fig. 7.
-Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (8).
-Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig.
+With the two conditions we can obtain the smooth radial wave function and the energy spectrum of the monopolium.
+The radial wave function for n=1,2,,5n=1,2,...,5 is shown in Fig.
+8.
+However, the condition (65) which determines the energy spectrum of the monopolium is not easy to solve, and it can only be solved numerically.
+Fortunately we could solve it, and the resulting energy spectrum is displayed in Fig.
 9 (for $l=0$).
+Figure 8: The radial monopolium wave function R(x) for $n=1,2,3,4,5$ shown in (62), (64), and (65).
 We see that as $x_{0}$ goes to zero, the binding energies approaches the values given by the atomic model
@@
 $\epsilon_{HO}=\frac{3\sqrt{x_{0}}-3-2l-4n}{x_{0}^{3/2}},~~~~n=0,1,\ldots$ (67)
-The value of $x_{0}$ we are interested is given by $r_{0}\sim 1/M_{W}$, or
+The value of $x_{0}$ we are interested in is given by $r_{0}\sim 1/M_{W}$, or
 $x_{0}=\frac{M}{2\alpha}r_{0}\simeq 7,767,173.$ (68)
-For such a high values of $x_{0}$, the binding energies are almost indistinguishable from binding energies of harmonic oscillator.
+For such a high values of $x_{0}$, the binding energies are almost indistinguishable from the binding energies of harmonic oscillator.
 This means that the ground state energy of the monopolium (for $l=0$) is roughly
@@
 Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium.
-In particular, this tells that the monopole-antimonopole pair with mass 22.0 TeV makes the monopolium bound state of mass of 5.2 TeV at the ground state.
+In particular, this tells that the monopole-antimonopole pair with mass 22 TeV makes the monopolium bound state of mass of 5.2 TeV at the ground state.
 This is precisely what we have hoped for.
+Figure 9: The energy spectrum $\sqrt{\epsilon}$ of the monopolium with $l=0$ fixed by (65), as function of $\sqrt{x_{0}}$.
+The horizontal dotted lines are the energy levels of the hydrogen atom given by (66), and the dashed lines represent harmonic oscillator energy levels given by (67).
 Furthermore, the Bohr radius of the monopolium, which can be read of from the exponential decay of $R_{\rm II}$ from (64) is given by
@@
 This confirms that the monopolium bound state with the modified potential (7) can give us the desired binding energy which could help LHC to produce the bound state even when the monopole mass becomes bigger than 7 TeV.
-Notice that in the above analysis we have assumed the monopole mass to be 11.0 TeV to have the monopolium mass 5.2 TeV.
+Notice that in the above analysis we have assumed the monopole mass to be 11 TeV to have the monopolium mass 5.2 TeV.
 And (69) tells that the monopolium mass becomes 3.3 TeV when the monopole mass becomes 7 TeV.
@@
 Despite the obvious difference between our potential (55) and this, we notice that the characteristic features of the two monopoliums are quite similar.
-In particular, in both cases the monopolium masses are slightly below the half of the monopole’s mass and the radiuses are of the order of $10^{-18}cm$.
+In particular, in both cases the monopolium masses are slightly below the half of the monopole mass and the radius are of the order of $10^{-18}cm$.
 This implies that our results discussed above are reliable.
 VII Discussion
-An urgent task at LHC is to discover the electroweak monopole.
+An urgent task at LHC is to discover the electroweak monopole which exists within the standard model.
 Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it.
 But the problem is that intuitively the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV.
-And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, LHC cannot produce it.
+And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, the present LHC cannot produce it.
 In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition [8, 20].
 In this paper we have shown that the question if the present LHC could produce the monopole or not critically depends on the monopole production mechanism at LHC.
-Our work in this paper tells the followings.
+Our work in this paper strongly implies the followings.
 First, if the monopole production mechanism at LHC is the topological thermal fluctuation of the Higgs vacuum, there is practically no hope that it could produce the monopole, simply because the temperature of the p-p fireball is too low to produce the monopole thermally.
+And this would be the case even with the new FCC.
 This is unexpected, but if the recent ALICE measurement of the temperature of the p-p fireball is trustable, this conclusion is unavoidable.
-Second, if the monopole production mechanism at LHC is the Drell-Yan process, the present 14 TeV LHC could produce the monopole if the monopole mass becomes less than 7 TeV.
-This was expected.
-What is new is that, more probably LHC could produce the monopole in the form of the monopolium bound states for any reasonable mass of the monopole, as far as the mass does not exceed 29,7 TeV.
+Second, if the monopole production mechanism at LHC is the Drell-Yan process, the present 14 TeV LHC could produce the monopole only if the monopole mass becomes less than 7 TeV. This was expected.
+In this paper we discussed a new monopole production mechanism, the monopole production in the form of the monopolium bound state, which could allow LHC to produce the electroweak monopole for any reasonable mass of the monopole, as far as the mass does not exceed 29,7 TeV.
 This is because the monopolium binding energy greatly reduces the mass of the monopolium.
-This tells that the present LHC could circumbent the energy constraint and produce the electroweak monopole even when the monopole mass becomes bigger than 7 TeV. This is remarkable.
+This tells that the present LHC could circumbent the energy constraint and produce the electroweak monopole in the form of the monopolium bound state, even when the monopole mass becomes bigger than 7 TeV.
+In this case we MoEDAL, ATLAS, and CMS could detect the decay modes of the monopolium and thus confirm the existence of the electroweak monopole at LHC.
+This is remarkable.
 Of course, with the Schwinger mechanism at LHC we can have the same conclusion, if the p-p fireball could create strong enough magnetic background which can actually produce the monopole-antimonopole pairs.
 But it is by no means clear how the p-p fireball at LHC could generate such a strong magnetic background.
-To summarize, our results tells that a most probable way for LHC to produce the monopole is the Drell-Yan process, and the 14 TeV LHC could more likely produce the monopole in the form of the monopolium bound states than as individual monopoles.
+To summarize, our results tells that a most probable way for LHC to produce the monopole is the monopole production in the form of the monopolium, and the 14 TeV LHC could more likely produce the monopole in the form of the monopolium bound states rather than as individual monopoles.
 In this case the most realistic way to detect the monopole at LHC is to detect the two photon decay modes of the monopoliums.
-And in principle ATLAS and CMS could do that.
-But there is the possibility that the LHC (with the next energy upgrading) may not produce the monopole signal at all.
-This could mean that the monopole production mechanism at LHC is probably not the Drell-Yan process.
+And in principle MoEDAL, ATLAS, and CMS could do that.
+Finally, it should also be emphasized that there is the possibility that the LHC may not produce the monopole signal at all.
+This is because neither the Drell-Yan process nor the Schwinger mechanism could turn out to be the monopole production mechanism at LHC.
 In this case the only way to detect the electroweak monopole could be to look for the remnant monopoles produced during the early universe, with the “cosmic” MoEDAL.
@@
 PB and FB are supported by the Institute of Experimental and Applied Physics, Czech Technical University in Prague and the Research Centre for Theoretical Physics and Astrophysics, Institute of Physics, Silesian University in Opava.
-YMC is supported in part by the National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University, Korea.
+YMC is supported in part by the National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999 and 2025-R1A2C17064731) and by Center for Quantum Spacetime, Sogang University, Korea.